### v2

# How much should we care about what others know? Jump signals in optimal investment under relative performance concerns

Abstract.
We present a multi-agent and mean-field formulation of a game between investors who receive private signals informing their investment decisions and who interact through relative performance concerns. A key tool in our model is a Poisson random measure which drives jumps in both market prices and signal processes and thus captures common and idiosyncratic noise. Upon receiving a jump signal, an investor evaluates not only the signal’s implications for stock price movements but also its implications for the signals received by her peers and for their subsequent investment decisions. A crucial aspect of this assessment is the distribution of investor types in the economy. These types determine their risk aversion, performance concerns, and the quality and quantity of their signals. We demonstrate how these factors are reflected in the corresponding HJB equations, characterizing an agent’s optimal response to her peers’ signal-based strategies. The existence of equilibria in both the multi-agent and mean-field game is established using Schauder’s Fixed Point Theorem under suitable conditions on investor characteristics, particularly their signal processes. Finally, we present numerical case studies that illustrate these equilibria from a financial-economic perspective. This allows us to address questions such as how much investors should care about the information known by their peers.
PB gratefully acknowledges financial support by Deutsche Forschungsgemeinschaft through CRC/TRR 388 “Rough Analysis, Stochastic Dynamics and Related Fields”, Projects B03 and B05. GS gratefully acknowledges financial support by Deutsche Forschungsgemeinschaft through IRTG 2544 “Stochastic Analysis in Interaction”.
Mathematical Subject Classification (2020):
91G10, 91A06, 91A16, 91B06
Keywords:
Optimal investment, jump signals, relative performance concern, idiosyncratic information
1. Introduction
Information is a key driver of investment decisions, especially when shaped by idiosyncratic signals about impending stock price shocks. When combined with relative performance concerns–—the competitive drive to measure one’s wealth against the average–—this creates a compelling dynamic: Investors must not only interpret their own signals but also account for their peers’ information flows and resulting actions, leading to questions such as: How much should we care about what others know?
In this paper, we propose a simple and tractable framework that allows us to address such questions. We combine the works of [BK22] and [LZ19] to obtain a financial market model featuring a heterogeneous population of agents who make stragegic decisions based on dynamically generated idiosyncratic signals on imminent price shocks. Their goal is to maximize the respective expected utility from a ratio between their terminal wealth and a measure of their peers’ average wealth.
We will start with a finite number of investors who trade between their individual risk-free bank accounts that pay a constant interest rate and individual risky assets modeled by exponential Lévy processes. These processes are driven by an idiosyncratic Brownian motion, a common Brownian motion and a Poisson random measure generating both common and idiosyncratic noise. Such an asset specialization reflects real-world observations: investors often concentrate on stocks they are familiar with (cf. [Mer87]), are geographically bound to (cf. [CM99]), or they specialize in certain industries (cf. [KSZ05]). That said, the possibility of one common stock market for all investors is a special case of this setting.
Additionally, our investors have access to individual signal processes which augment their classical predictable information flow to Meyer-measurable strategies ([Len80, BB19]). So, each investor’s information flow is modeled by a $\sigma$-field which also includes information from their individual signal process and, hence, allows for a proactive response upon signal reception. In [BK22], the authors develop a rigorous theory for these kinds of jump signals and signal-driven strategies in the context of a single investor’s optimal investment problem. Importantly, in the present paper, the information provided by signals may differ between investors, for example, some investors may receive more frequent, yet less reliable signals than other investors.
To capture investor heterogeneity, we not only consider differences in the information provided by signals but also in risk aversion and performance concern parameters. Empirical evidence from [Koi14] supports the presence of heterogeneity in risk aversion among fund managers and gives good reason to consider investors that differ in their risk aversion and their performance concerns. Each investor’s characteristics are collected in her type (vector), whose distribution is assuemd to be common knowledge. This assumption is (at least partially) justified by the fact that investors often post their returns publicly, such that, for example, their risk preferences can be inferred from this information (cf. [Koi14]). Such transparency is valuable for investors when deciding their strategies, particularly in the presence of interaction.
Strategic interactions among investors arise from various motivations. For individual investors, these interactions may stem from aspirations to accumulate more wealth than their peers, cf. for example [Abe90, DKK08]. For fund managers, interactions may reflect career incentives or efforts to attract greater inflows from clients, cf. [BM15]. We model these interactions through relative performance concerns, which balance the trade-off between absolute and relative wealth. Indeed, our investors adopt power utility where a concern parameter mediates the impact from terminal wealth and their peers’ geometric mean terminal wealth. Optimal investment under such relative performance concerns has been studied in [FdR11], [ET15], [BM15], [BLdR17], and [LZ19], to name a few.
The relative performance criteria lead to an interdependence of the investors’ target functionals and thus introduce a stochastic game of optimal investment which calls for the search for Nash equilibria. To address this, we focus on a novel class of signal-driven strategies, where investment decisions depend only on the signal (if any) received at each time. The existence of a Nash equilibrium is established in two steps. First, we solve a single investor’s best response problem when her peers follow fixed signal-driven strategies. Using the approach in [BK22], we formulate an HJB equation and prove the corresponding verification results, which characterizes best response strategies as solutions to one-dimensional concave optimization problems and identifies them to also be signal-driven. This allows us to introduce, in a second step, a fixed-point mapping that encodes the so-called consistency condition. This condition ensures that the best response controls align with the fixed environment of peers’ strategies, meaning each strategy equals its own best response strategy. Under suitable assumptions on the investors’ types, we use Schauder’s fixed point theorem to establish existence of an equilibrium.
The focus on signal-driven strategies is even more natural in a mean field setting where a single agent’s action does not impact the overall average considered by her peers. We thus provide a formulation for the corresponding mean field game via randomization over the investors’ type vectors. The random type then represents the representative investor for which we can again solve the best response problem having fixed the competitive environment. An important intermediate step is the derivation of the geometric mean wealth in the mean field setting by conditioning on the common noise. For this, we will carefully disentangle what part of the randomness is common to all investors and which part is idionsyncratic (and thus averaged out in the mean field).
Finally, we prove the existence of a mean field equilibrium in two different scenarios. First, we analyze the case of finitely many investor types under suitable assumptions, utilizing arguments analogous to those employed in the multi-agent game. Secondly, when the mean field is allowing for infinitely many different types, we device an alternative fixed-point map which allows us to prove existence of an equilibrium provided there are only finitely many distinct common shocks.
In numerical experiments, we investigate how the presence of idiosyncratic investment signals impacts investors with relative performance concerns. For this, we fully specificy a mean field game with just two types of investors and introduce a suitable kind of certainty equivalent for one of these types. By varying parameters of our types, this allows us to address questions such as how much an investor should care about what is known by her peers. Our findings reveal a dual effect: investors who value relative wealth prefer their peers to have strong, and even superior, signal quality as better-informed peers help mitigate relative wealth fluctuations. Additionally, they support increased competitiveness among their peers, as this aligns ambitions to stabilize fluctuations in relative wealth. However, investors also favor a relatively uniform signal quality across the population. Indeed, how much we should care about others’ knowledge turns out to depend on the reliability of their knowledge compared to our own: If it is equally reliable, its impact is minimal; if it differs, it becomes significant, regardless of whether it is more or less reliable.
Related Literature. Multi-agent and mean field games provide a powerful framework for studying strategic interactions among a large number of agents, such as investors in financial markets. While multi-agent games capture the finite-population dynamics where individual decisions indirectly or directly affect others, mean field games approximate these interactions in large populations by considering the aggregate behavior, significantly simplifying the mathematical analysis and computational efforts. Mean field games, introduced independently by [HMC06] and [LL07], have gained popularity in financial mathematics over the last two decades. It is beyond the scope of this paper to provide a comprehensive introduction to mean field games, and for this, we refer to [BFY13], [CDLL19] and [CD18]. For our context, we would like to direct the attention to [ET15], [BLdR17] and [LZ19]. The former two both consider optimal investment for exponential utility with relative performance concerns. In [ET15], the authors establish in a Black-Scholes market the existence and uniqueness of a Nash equilibrium for both unconstrained agents and those subject to portfolio constraints. Similarly, [BLdR17] investigates both an $n$-agent and a mean field game framework and employs coupled quadratic backward stochastic differential equations (BSDEs) to study equilibrium pricing under social interaction. Finally, the latter considers for CARA and CRRA preferences a multi-agent and a mean field game with relative performance concerns in a Black-Scholes market with common noise Brownian motion. The authors obtain unique time-independent constant equilibria. In particular, [LZ19] was followed by much subsequent work in various directions including [BG23], [BWY24a], [BWY24b], [LS19], [dRP22] and [FZ23].
None of these works, however, considers the interplay of asymmetric information about market jumps in the context of relative utility maximization. Indeed, there is not much literature treating jumps as both part of the common and idiosyncratic noise. Concerning this, [BH24] considers a new type of mean-field portfolio games which exhibits both idiosyncratic and common noise components in terms of jumps through integer-valued random measures. The authors in [BH24] characterize equilibria for mean-field portfolio games of both optimal investment and hedging under relative performance concerns with respect to exponential (CARA) utility preferences by solutions to suitable McKean-Vlasov FBSDEs with jumps, for which they show existence and uniqueness.
We take a different path and consider the Poisson random measure to place at common jump times both idiosyncratic and common jump marks. These jump marks account both for jumps in the stock prices and for idiosyncratic signals which may transport (imperfect) information about the price jumps. In particular, our investors can be viewed as ‘small insiders’ who do not affect the price and for whom any possible information advantage on imminent price shocks is short-lived. In fact, it is this short-term information advantage which affords us a lot of mathematical tractability as it precludes highly complex strategic considerations such as concern about information leakage (What do my actions betray about my private knowledge?) or bluffing (Is my counterpart really thinking what I expect her to?) which make other games with private information flows so difficult to analyze.
The remainder of this paper is organized as follows: Section 2 first introduces the multi-agent game, discusses then the best-response problem and concludes with proving existence of Nash equilibria. Section 3 shows how to follow this approach in a mean field setting. Finally, Section 4 illustrates numerical experiments and addresses some financial-economic questions.
2. The Multi-Agent Game
This section introduces a multi-agent framework where investors navigate interactive utility maximization in the presence of jump signals. Subsection 2.1 presents the market model, the investors’ signal processes and their sets of admissible strategies. Subsection 2.2 specifies the individual utility functions, accounting for relative performance concerns and heterogeneity in risk aversion and concern parameters, which shape the investors’ best response problems. Section 2.3 provides a best response map for a single investor interacting with investors who employ fixed signal-driven strategies. Subsection 2.4 provides a proof of existence of a Nash equilibrium. Finally, Subsection 2.5 establishes a convergence result for aggregate wealth as the number of investors $n$ goes to $\infty$, paving the way for the mean field game.
2.1. The Financial Market Model with Jump Signals
In the following, we introduce the market model and specify the investors’ admissible strategies.
Let $T>0$ be a finite time horizon and $n\in\mathbb{N}_{\geq 1}$. In our model, randomness is introduced via a probability space $(\Omega,\mathfrak{A},\mathbb{P})$ which supports $n+2$ independent standard Brownian motions $W^{0},W^{1},...,W^{n+1}$ and an independent Poisson random measure $N=N(dt,de)$. The Poisson measure is compensated by the intensity measure $dt\otimes\nu(de)$ for a positive finite measure $\nu$ on some Polish mark space $E$ equipped with the corresponding Borel-$\sigma$-field $\mathcal{E}=\mathcal{B}(E)$. We denote by $\mathcal{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ the right-continuous filtration induced by the completions of
$\sigma((W^{0}_{s},W^{i}_{s},N([0,s]\times A)):s\in[0,t],A\in\mathcal{E},i=1,...,n+1),\quad t\in[0,T],$
and by $\mathcal{P}=\mathcal{P}(\mathcal{F})$ the predictable $\sigma$-field generated by all continuous $\mathcal{F}$-adapted processes on $\Omega\times\mathbb{R}_{+}$.
Market Model. We consider a finite number of $n+1$ investors that trade between their individual risk-free bank accounts $S^{0i}$ with interest rate $r^{i}\in\mathbb{R}$ and an individual risky stocks $S^{i}$. The asset price dynamics are given by
(1) $\begin{split}S_{0}^{0i}=s^{0i}_{0}>0,\quad dS_{t}^{0i}&=r^{i}S_{t-}^{0i}dt,\quad t\in[0,T],\\ S^{i}_{0}=s_{0}^{i}>0,\quad\frac{dS^{i}_{t}}{S^{i}_{t-}}&=\kappa^{i}dt+\sigma^{i}dW^{i}_{t}+\sigma^{0i}dW^{0}_{t}+\int_{E}\eta^{i}(e)N(dt,de),\quad t\in[0,T],\end{split}$
for constants $\kappa^{i}\in\mathbb{R}$ and $\sigma^{i},\sigma^{0i}\geq 0$ such that $\sigma^{i}+\sigma^{0i}>0$ for $i=1,...,n+1$ and for measurable functions $\eta^{i}:E\to\mathbb{R}$ for $i=1,...,n+1$. We will assume
(2) $\nu(\{\eta^{i}<-1\})=0,$
to guarantee that each $S^{i}$ remains nonnegative at all times.
So, while the Brownian motions $W^{i}$ are idiosyncratic noise for respective investor $i=1,...,n+1$, the Brownian motion $W^{0}$ accounts for common noise. The Poisson random measure $N$ intoduces jumps by placing marks $e\in E$ at common times for all investors. These marks translate into jump sizes $\eta^{i}(e)$ which may differ between the investors $i=1,\dots,n+1$. The case closest to reality where all investors trade between the same common bond and the same common stock corresponds to the parameter choice where, for all $i=1,...,n+1$, $r^{i}=r$, $\sigma^{i}=0$, $\kappa^{i}=\kappa$, $\sigma^{0i}=\sigma^{0}$ and $\eta^{i}(e)=\eta(e)$ for some $r\in\mathbb{R}$, $\kappa\in\mathbb{R}$, $\sigma^{0}>0$ and some measurable function $\eta:E\to\mathbb{R}$. It is merely for notational simplicity that we refrain from specifying a multivariate model with several stocks. We refer to [ET15], [FdR11] and [LZ19] for an in-depth study of the single stock case, albeit without jumps.
Wealth Dynamics. Let $x^{i}_{0}>0$ be the $i$th investor’s initial wealth. The $i$th investor’s investment strategies are described by a stochastic process $\tilde{\phi}^{i}=(\tilde{\phi}^{i}_{t})_{t\in[0,T]}$ specifying the fraction of wealth invested in the risky asset $S^{i}$ at times $t\in[0,T]$. So, the investor’s wealth process $X^{i}:=(X^{i}_{t})_{t\in[0,T]}:=(X^{\tilde{\phi}^{i}}_{t})_{t\in[0,T]}$ will evolve accordingly as
(3) $\begin{split}X_{0}^{i}=x^{i}_{0},\quad\frac{dX_{t}^{i}}{X^{i}_{t-}}=(1-\tilde{\phi}^{i}_{t})\frac{dS_{t}^{0i}}{S^{0i}_{t-}}+\tilde{\phi}^{i}_{t}\frac{dS_{t}^{i}}{S^{i}_{t-}},\quad t\in[0,T].\end{split}$
Note that, as the jump intensity is finite, this will be well-defined for any real-valued predictable $\tilde{\phi}^{i}$ satisfying
[…]
In fact, since we assumed the intensity measure $\nu$ to be finite, there will be only finitely many jumps in asset prices in each scenario so that wealth dynamics will even be defined for optimal $\tilde{\phi}^{i}$ with the above square-integrability. We refer to [BK22] for the additional integrability required in the case where $\nu$ allows infinitely many jumps and refrain from considering this possibility here for ease of exposition.
Jump Signals via Meyer-$\sigma$-Fields. As already mentioned, when deciding over the proportions invested in the risky asset, investors may, at times, have the possibility to recourse to some signal that gives them extra information on impending jumps in the stock price.
The signal process of the $i$th investor is defined as
(4) $Z^{i}_{t}=\int_{\{t\}\times E}\zeta^{i}(e)N(ds,de),\quad t\in[0,T],$
for a measurable function $\zeta^{i}:E\to\mathbb{R}$. Observing a value $Z^{i}_{t}\not=0$ allows the investor to infer (typically imperfect) information about the mark $e\in E$ that has been set by the Poisson random measure. As a consequence, she will revise the probability of the corresponding asset prices jumps and even start musing who among her fellow investors will also have received a signal, what this signal may have been and what it reaction it will trigger. Obviously, we have to allow our investor to account for the received signal in her investment choice and so we let her choose a stategy that is measurable with respect to the Meyer-$\sigma$-field
$\Lambda^{i}:=\mathcal{P}\vee\sigma(Z^{i}).$
We refer to [Len80], [Kar81] and [BB19] for the detailed theory of Meyer-$\sigma$-fields. For our purposes, it suffices to note how this is reflected in the mathematical structure of the investors’ investment strategies:
Lemma 2.1 (cf. [BK22], Corollary 2.2).
A real-valued process $\tilde{\phi}^{i}$ is measurable with respect to the Meyer-$\sigma$-field $\Lambda^{i}=\mathcal{P}(\mathcal{F})\vee\sigma(Z^{i})$ if and only if it can be written in the form
$\tilde{\phi}^{i}_{t}(\omega)=\hat{\phi}^{i}_{t}(\omega,Z^{i}_{t}(\omega)),\quad(\omega,t)\in\Omega\times[0,T],$
for some $\mathcal{P}(\mathcal{F})\otimes\mathcal{B}(\mathbb{R})$-measurable field $\hat{\phi}^{i}:[0,T]\times\Omega\times\mathbb{R}\to\mathbb{R}$.
In light of Lemma 2.1, investor $i$ has to decide in a predictable manner about all the respective positions $\hat{\phi}^{i}_{t}(\omega,z)$, $z\in\mathbb{R}$, she will take if in the next moment she was to receive a specific signal $z=Z^{i}_{t}(\omega)$.
Admissible strategies. It will be crucial to understand what information each investor can derive from her signals. For instance, the kind of signals received (or not) at any one time will affect the investors’ positions that keep their respective wealth nonnegative almost surely.
We follow the disintegration argument in [BK22, pp.1306-1307]. That is, we write
$\nu(de\cap\{\zeta^{i}\neq 0\})$ $=\int_{\zeta^{i}(E)\setminus\{0\}}K^{i}(z,de)\mu^{i}(dz)\quad\text{ for }\mu^{i}:=\nu\circ(\zeta^{i})^{-1}\text{ on }\mathbb{R}.$
Note that this disintegration is indeed possible since $\nu$ is finite and positive, cf. [JS03, II. 1.2, p. 65]. The finite measure $\mu^{i}(dz)$ describes the frequency with which the $i$th investor receives a non-zero signal $z\in\zeta^{i}(E)\setminus\{0\}$, while $K^{i}(z,de)$ describes the a posteriori distribution of the marks $e\in E$ given a signal $z\neq 0$. Note that $K^{i}(z,\{\zeta^{i}=z\})=K^{i}(z,E)=1$.
In moments when investor $i$ receives a signal $z\neq 0$, she can adopt any position $\phi^{i}_{t}(z)$ that guarantees nonnegative wealth after any imminent price jump $\eta^{i}(e)$ that may come from any mark $e$ that is compatible with the observed signal in the sense that $\zeta^{i}(e)=z$. To formalize this, we introduce tight jump bounds
$\underline{\eta}^{i}(z):=K^{i}(z,\cdot)\text{-ess inf }\eta^{i}\geq-1\text{ and }\bar{\eta}^{i}(z):=K^{i}(z,\cdot)\text{-ess sup }\eta^{i}$
which allow us to describe the positions ensuring nonnegative wealth by the interval
$\Phi^{i}(z):=[-1/\bar{\eta}^{i}(z),-1/\underline{\eta}^{i}(z)],\quad z\in\zeta^{i}(E)\setminus\{0\}.$
Note that these intervals are compact under our henceforth standing no arbitrage condition
(5) $\underline{\eta}^{i}(z)<0<\bar{\eta}^{i}(z)\text{ for $\mu^{i}$-a.e. }z\in\zeta^{i}(E)\setminus\{0\},$
which rules out that investors learn for sure the direction of the next price jump given a signal $z\neq 0$.
In case of no signal, i.e. when $Z^{i}_{t}=z=0$, the investor learns that (if any) a mark can only be set in $\{\zeta^{i}=0\}$. Thus, if there are no unsignaled shocks, i.e. if $\nu(\zeta^{i}=0,\eta^{i}\neq 0)=0$, she can choose any position in $\Phi^{i}(0):=\mathbb{R}$. If, by contrast, there may be unsignaled shocks, i.e. if $\nu(\zeta^{i}=0,\eta^{i}\neq 0)>0$, we ensure nonnegative wealth by restricting to positions in
$\Phi^{i}(0):=\begin{cases}[-1/\bar{\eta}^{i}(0),-1/\underline{\eta}^{i}(0)],\quad&\text{ if }\underline{\eta}^{i}(0)<0<\bar{\eta}^{i}(0),\\ [-1/\bar{\eta}^{i}(0),+\infty),\quad&\text{ if }0\leq\underline{\eta}^{i}(0)\leq\bar{\eta}^{i}(0),\\ (-\infty,-1/\underline{\eta}^{i}(0)],\quad&\text{ if }\underline{\eta}^{i}(0)\leq\bar{\eta}^{i}(0)\leq 0,\end{cases}$
where the jump bounds in this case are
$\underline{\eta}^{i}(0)$ $:=\nu(\cdot\cap\{\zeta^{i}=0,\eta^{i}\neq 0\})\text{-ess inf }\eta^{i}\geq-1,$
$\bar{\eta}^{i}(0)$ $:=\nu(\cdot\cap\{\zeta^{i}=0,\eta^{i}\neq 0\})\text{-ess sup }\eta^{i}.$
Notice that the condition $\sigma^{i}+\sigma^{0i}>0$ ensures no arbitrage also when unsignaled shocks can only go in one direction since it ensures that at least one of the Brownian motions $W^{i}$ or $W^{0}$ keeps stock prices fluctuating suitably.
This allows us to fix the set of admissible strategies which keep our investors’ wealth levels even strictly above zero as follows:
Proposition 2.2 (cf. [BK22], Corollary 2.5).
Suppose (2) and (5) hold. Let, for all $z\in\zeta^{i}(E)\cup\{0\}$, $\bar{\Phi}^{i}(z)$ be a compact interval included in $\{0\}\cup\operatorname*{int}\Phi^{i}(z)$. Then, any $\Lambda^{i}$-measurable strategy $\tilde{\phi}^{i}=\hat{\phi}^{i}(Z^{i})$ in the set of admissible strategies
[…]
admits wealth dynamics $X^{i}=X^{\tilde{\phi}^{i}}=X^{\hat{\phi}^{i}}$ solving (3) which remain strictly positive throughout $[0,T]$ almost surely.
With the financial market model, the signal and wealth processes of investors subject to admissible strategies at hand, we can move on to their individual optimal investment problems.
2.2. Risk Preferences, Relative Performance Concerns and Investor Types
We assume that, each investor $i=1,...,n+1$ has an individual constant relative risk aversion $\alpha^{i}\in(0,1)\cup(1,\infty)$. She assesses her wealth $x$ not only by itself, but also in view of her peers’ average wealth $\bar{x}$. Her specific concern for the others’ performance is measured by a parameter $\theta^{i}\in[0,1]$ which enters her utility function via
(6) $u_{i}(x,\bar{x}):=u_{\alpha^{i}}(x\bar{x}^{-\theta^{i}}):=\frac{(x\bar{x}^{-\theta^{i}})^{1-\alpha^{i}}}{1-\alpha^{i}}\quad\text{ for }x,\bar{x}>0.$
We furthermore let $u_{i}(0,0):=0$ and, for $x,\bar{x}\in(0,\infty)$, we fix
(7) $u_{i}(x,0):=\begin{cases}+\infty&\text{ if }\alpha^{i}\in(0,1),\\ 0&\text{ if }\alpha^{i}>1,\end{cases}\quad\text{ and }\quad u_{i}(0,\bar{x}):=\begin{cases}0&\text{ if }\alpha^{i}\in(0,1),\\ -\infty&\text{ if }\alpha^{i}>1.\end{cases}$
For investor $i$, her peers’ geometric average wealth at any time $t\in[0,T]$ is of course
(8) $\bar{X}_{t}^{-i}:=\left(\prod_{j\neq i}X^{i}_{t}\right)^{\frac{1}{n}},$
and so, comparing the terminal average $\bar{X}^{-i}_{T}$ with her own terminal wealth $X^{i}_{T}$, she will assess this investment outcome by its utility
(9) $u_{i}(X^{i}_{T},\bar{X}^{-i}_{T})=u_{\alpha^{i}}(X^{i}_{T}/(\bar{X}_{T}^{-i})^{\theta^{i}})=\frac{1}{1-\alpha^{i}}\bigg{(}\frac{X^{i}_{T}}{(\bar{X}_{T}^{-i})^{\theta^{i}}}\bigg{)}^{1-\alpha^{i}}.$
Given this utility structure, a concern parameter $\theta^{i}$ close to 00 corresponds to low (or even no) relative performance concerns while $\theta^{i}$ close to $1$ corresponds to high relative performance concerns of the $i$th investor. In other words, the higher the value of $\theta^{i}$, the more the investor’s concern shifts from her networth $X^{i}_{T}$ to its size compared to the average wealth $\bar{X}^{-i}_{T}$. Clearly, the geometric mean is chosen for tractability, as it allows us to exploit the homogeneity of the utility function. Note also that degeneracies will occur when one of the investors has $X^{j}_{T}=0$ as this entails $\bar{X}^{-i}_{T}=0$, a singularity we will be able to avoid due to our notion of admissible strategies, see Proposition 2.2.
Remark 2.3.
We model the interaction and relative performance concerns as in [LZ19, Sec. 3]. The authors in [LZ19] investigate both constant absolute risk aversion (CARA) and constant relative risk aversion (CRRA) utility separately. Since [BK22] considers CRRA utility only, we focus here excusively on such utility functions as well, making this paper correspond to the special case where $\theta^{i}=0$. Note also that we omit the case $\alpha^{i}=1$ corresponding to log-utility, because investors with log-utility are always indifferent to their peers, regardeless of $\theta^{i}$.
The characteristics of each investor $i=1,...,n+1$ are now fully described and can be summerized by her corresponding type (vector)
$\mathtt{t}^{i}=(x_{0}^{i},r^{i},\kappa^{i},\sigma^{i},\sigma^{0i},\eta^{i},\zeta^{i},\alpha^{i},\theta^{i})$
from the type space
$\mathtt{T}:=(0,\infty)\times\mathbb{R}^{2}\times([0,\infty)^{2}\setminus\{0,0\})\times L^{0}(E)\times L^{0}(E)\times((0,\infty)\setminus\{1\})\times[0,1]$
where $L^{0}(E)$ denotes the set of measurable functions mapping $E$ to $\mathbb{R}$.
Noting that a collection of type vectors $(\mathtt{t}^{1},\ldots,\mathtt{t}^{n+1})$ captures all relevant information about the investors involved, we henceforth assume types to be common knowledge, that is all investors know their own type and the other investors’ types.
It will be convenient for our analysis to make the following standing assumption:
Assumption 2.4.
Our investors’ types $(\mathtt{t}^{1},...,\mathtt{t}^{n+1})\in\mathtt{T}^{n+1}$ are such that
  1. $(\mathtt{t}_{1})$
    the stock price remains nonnegative, i.e. $\nu(\{\eta^{i}<-1\})=0$,
  2. $(\mathtt{t}_{2})$
    the no-arbitrage condition holds, i.e. $\underline{\eta}^{i}(z)<0<\bar{\eta}^{i}(z)\text{ for $\mu^{i}$-a.e. }z\in\zeta^{i}(E)\setminus\{0\}$,
  3. $(\mathtt{t}_{3})$
    the signal map $e\mapsto\zeta^{i}(e)$ takes only finitely many values, and
  4. $(\mathtt{t}_{4})$
    jump sizes are sufficiently integrable, in the sense that
    $\int_{E}(1+\eta^{i}(e))^{p}\nu(de)<\infty\text{ for all }p>0.$
Here, $(\mathtt{t}_{3})$ means that, while there may be an infinite range of possible marks $e\in E$ placed by the Poisson random measure, each signal function $\zeta^{i}$, $i=1,\dots,n+1$ classifies this information into a range of only finitely many discrete categories. This is not unreasonable in a real-world context, as this classification helps reduce the complexity of decision making by providing a manageable amount of information that is still sufficiently useful. Mathematically, $(\mathtt{t}_{3})$ will be crucial to get the necessary continuity and compactness we need for our proof of existence of a Nash equilibrium. Condition $(\mathtt{t}_{4})$ is a non-degeneracy assumption which ensures that expected utility functions remain finite.
Moreover, this assumption allows to show a very convenient property:
Lemma 2.5.
Following an admissible strategy $\tilde{\phi}^{i}=\hat{\phi}^{i}(Z^{i})\in\mathcal{A}^{i}$, each investor $i=1,\dots,n+1$ is uniformly shielded from ruin in the sense that her investment returns satisfy
$0<\nu\text{-}\operatorname*{ess\,inf}(1+(\hat{\phi}^{i}\circ\zeta^{i})\eta^{i}).$
Proof.
By $(\mathtt{t}_{3})$, $\zeta^{i}(E)\cup\{0\}$ is a finite set. Thus, we can minimize over all possible signals $z$, positions $x$ and jumps $y$ to find that
$1+(\hat{\phi}^{i}_{t}\circ\zeta^{i})\eta^{i}\geq\min_{z\in\zeta^{i}(E)\cup\{0\}}\min_{x\in\bar{\Phi}^{i}(z)}\min_{y\in[\underline{\eta}^{i}(z),\bar{\eta}^{i}(z)]}1+xy.$
Further, our restriction to positions in $\bar{\Phi}^{i}(z)$ ensures that $-1\notin\{xy:x\in\bar{\Phi}^{i}(z),y\in[\underline{\eta}^{i}(z),\bar{\eta}^{i}(z)]\}$, for respective signals $z\in\zeta^{i}(E)\cup\{0\}$, and thus the right-hand side is strictly greater than zero. ∎
Now, given the other investors’ strategies $\tilde{\phi}^{-i}:=(\tilde{\phi}^{1},...,\tilde{\phi}^{i-1},\tilde{\phi}^{i+1},...,\tilde{\phi}^{n+1})$ and their corresponding average wealth $\bar{X}^{-i}_{T}$ of (8), the $i$th investor’s best response problem is to find a strategy $\tilde{\phi}^{i}$ which
(10) $\text{maximizes }J^{i}(\tilde{\phi}^{i};\tilde{\phi}^{-i}):=\mathbb{E}\big{[}u_{i}(X_{T}^{i},\bar{X}_{T}^{-i})\big{]}\text{ over }\tilde{\phi}^{i}=\hat{\phi}^{i}(Z)\in\mathcal{A}^{i},$
subject to her wealth dynamics
(11) $\begin{split}X^{i}_{0}&=x_{0}^{i},\\ \frac{dX^{i}_{t}}{X_{t-}^{i}}&=(r^{i}+\hat{\phi}_{t}^{i}(0)(\kappa^{i}-r^{i}))dt+\sigma^{i}\hat{\phi}_{t}^{i}(0)dW^{i}_{t}+\sigma^{0i}\hat{\phi}_{t}^{i}(0)dW^{0}_{t}+\int_{E}\eta^{i}(e)\hat{\phi}_{t}^{i}(\zeta^{i}(e))N(dt,de).\end{split}$
Recognizing the interdependence of the investors’ optimization problems through the geometric mean, we find that we are dealing with an $n$-dimensional stochastic differential game which is best understood in terms of its Nash equilibria:
Definition 2.6.
A vector of admissible strategies $\tilde{\pi}^{\star}=(\tilde{\pi}^{1,\star},....,\tilde{\pi}^{n+1,\star})\in\mathcal{A}^{1}\times...\times\mathcal{A}^{n+1}$ is called a Nash equilibrium, if for each $i\in\{1,...,n+1\}$ we have
(12) $J^{i}(\tilde{\pi}^{i},\tilde{\pi}^{-i,\star})\leq J^{i}(\tilde{\pi}^{i,\star},\tilde{\pi}^{-i,\star})\text{ for all $\tilde{\pi}^{i}\in\mathcal{A}^{i}$}.$
So, as usual, a collection of strategies $\tilde{\pi}^{\star}=(\tilde{\pi}^{1,\star},....,\tilde{\pi}^{n+1,\star})\in\mathcal{A}^{1}\times...\times\mathcal{A}^{n+1}$ is called a Nash equilibirum if it is its own best response in the sense that each $\tilde{\pi}^{i,\star}$, $i=1,\dots,n+1$ solves (10) for $\tilde{\phi}^{-i}:=\tilde{\pi}^{-i,\star}$. While finding such best responses for arbitrary peers’ strategies $\tilde{\phi}^{-i}=\tilde{\pi}^{-i,\star}$ is a daunting task in general, it will turn out to be entirely feasible if they come from the following restricted class:
Definition 2.7.
Strategies $(\tilde{\phi}^{1},\dots,\tilde{\phi}^{n+1})=(\hat{\phi}^{1}(Z^{1}),\dots,\tilde{\phi}^{n+1})\in\mathcal{A}^{1}\times\dots\mathcal{A}^{n+1}$ are called (purely) signal-driven if each $\tilde{\phi}^{i}$, $i=1,\dots,n+1$, can be written in the form
(13) $\tilde{\phi}^{i}_{t}(\omega)=\phi^{i}(Z^{i}_{t}(\omega)),\quad(\omega,t)\in\Omega\times[0,T],$
for some $\mathcal{B}(\mathbb{R})$-measurable time-independent and deterministic function $\phi^{i}:\mathbb{R}\to\mathbb{R}$. We denote by
[…]
the set of admissible signal-driven strategies for the $i$th investor.
Definition 2.8.
A Nash equilibrium $\pi^{\star}=(\pi^{1,\star},....,\pi^{n+1,\star})$ is called a signal-driven Nash equilibrium if it is signal-driven in the sense of Definition 2.7, i.e. if $(\pi^{1,\star},....,\pi^{n+1,\star})\in\mathcal{A}_{\text{sig}}^{1}\times...\times\mathcal{A}_{\text{sig}}^{n+1}$.
Note that either notion of Nash equilibrium presented here is of open-loop type, that is, the equilibrium strategies are best responses within a class of strategies written in terms of the system’s driving noise.
2.3. HJB equation and Best Response Map for The Multi-Agent Game
The main result of this section is a verification theorem for best response controls for an investor who interacts with investors employing signal-driven strategies. As we will see, the best response in the set of all admissible strategies $\mathcal{A}^{i}$ is itself signal-driven, i.e., it comes from the smaller set $\mathcal{A}^{i}_{\text{sig}}$.
This result will be obtained by dynamic programming techniques similar to the case with one investor solved in [BK22]. As in [LZ19], the key difference here is of course that each investor $i\in\{1,...,n+1\}$ has to take into account her peers’ average mean wealth (8). When they persue admissible, signal driven strategies $\pi^{j}\in\mathcal{A}^{j}_{\text{sig}}$, $j\in\{1,...,n+1\}\setminus\{i\}$, we can use Itō’s formula to compute the dynamics of this average, which, conveniently, turns out to be an exponential Lévy process:
(14) $\begin{split}\bar{X}^{-i}_{t}=\bar{x}_{0}\exp\bigg{(}(\overline{\tau\pi})^{-i}t&+\frac{1}{n}\sum_{j\neq i}\sigma^{j}\pi^{j}(0)W^{j}_{t}+(\overline{\sigma^{0}\pi})^{-i}W^{0}_{t}\\ &+\int_{[0,t]\times E}\log((\overline{1+(\pi\circ\zeta)\eta})^{-i}(e))N(ds,de)\bigg{)}\end{split}$
with dynamics given by
(15) $\begin{split}\bar{X}^{-i}_{0}&=\bar{x}_{0}:=\bigg{(}\prod_{j\neq i}x^{j}_{0}\bigg{)}^{\frac{1}{n}}\\ \frac{d\bar{X}^{-i}_{t}}{\bar{X}^{-i}_{t-}}&=(\overline{\tau\pi})^{-i}dt+\frac{1}{2}[((\overline{\sigma^{0}\pi})^{-i})^{2}+(\overline{\sigma^{2}\pi^{2}})^{-i}]dt+\frac{1}{n}\sum_{j\neq i}\sigma^{j}\pi^{j}(0)dW^{j}_{t}\\ &\qquad\qquad\qquad+(\overline{\sigma^{0}\pi})^{-i}dW^{0}_{t}+\int_{E}\big{[}(\overline{1+(\pi\circ\zeta)\eta})^{-i}(e)-1\big{]}N(dt,de),\quad t\in[0,T],\end{split}$
where
$(\overline{\tau\pi})^{-i}$ $:=\frac{1}{n}\sum_{j\neq i}\big{[}r^{j}+\pi^{j}(0)(\kappa^{j}-r^{j})-\frac{1}{2}((\sigma^{j})^{2}+(\sigma^{0j})^{2})(\pi^{j}(0))^{2}\big{]},\quad(\overline{\sigma^{0}\pi})^{-i}:=\frac{1}{n}\sum_{j\neq i}\sigma^{0j}\pi^{j}(0)$
$(\overline{\sigma^{2}\pi^{2}})^{-i}$ $:=\frac{1}{n^{2}}\sum_{j\neq i}(\sigma^{j})^{2}(\pi^{j}(0))^{2}\text{ and }(\overline{1+(\pi\circ\zeta)\eta})^{-i}(e):=\prod_{j\neq i}(1+\pi^{j}(\zeta^{j}(e))\eta^{j}(e))^{\frac{1}{n}}.$
Bringing together the perspectives from [LZ19] and [BK22, pp.1312-1314], we can now derive the HJB equation for the best response problem (10). The corresponding value function is
$v^{i}(x^{i}_{0},\bar{x}_{0},T):=\sup_{\phi^{i}\in\mathcal{A}^{i}}\mathbb{E}\big{[}u_{i}(X^{i}_{T},\bar{X}^{-i}_{T})\big{]}.$
Assuming $v^{i}$ to be smooth, we expect by Itō’s formula that, for any admissible strategy $\tilde{\phi}^{i}=\hat{\phi}^{i}(Z)$, the value process $V_{t}^{i}:=V_{t}^{\phi^{i},\pi^{-i}}:=v^{i}(X^{i}_{t},\bar{X}^{-i}_{t},T-t)$, $t\in[0,T]$, to have the dynamics:
$v^{i}(X^{i}_{t},\bar{X}^{-i}_{t},T-t)$ $=v^{i}(x^{i}_{0},\overline{x}_{0},T)-\int_{0}^{t}\partial_{T}v^{i}(X^{i}_{s-},\bar{X}^{-i}_{s-},T-s)ds$
$\quad+\int_{0}^{t}\partial_{x}v^{i}(X^{i}_{s-},\bar{X}^{-i}_{s-},T-s)dX_{s}^{i,c}+\int_{0}^{t}\partial_{\bar{x}}v^{i}(X^{i}_{s-},\bar{X}^{-i}_{s-},T-s)d\bar{X}^{-i,c}_{s}$
$\quad+\int_{0}^{t}\partial_{x\bar{x}}v^{i}(X^{i}_{s-},\bar{X}^{-i}_{s-},T-s)d[X^{i},\bar{X}^{-i}]^{c}_{s}+\frac{1}{2}\int_{0}^{t}\partial_{xx}v^{i}(X^{i}_{s-},\bar{X}^{-i}_{s-},T-s)d[X^{i},X^{i}]_{s}^{\mathtt{c}}$
$\quad+\frac{1}{2}\int_{0}^{t}\partial_{\bar{x}\bar{x}}v^{i}(X^{i}_{s-},\bar{X}^{-i}_{s-},T-s)d[\bar{X}^{-i},\bar{X}^{-i}]_{s}^{c}$
$\quad+\sum_{0<s\leq t}\left[v^{i}(X^{i}_{s},\bar{X}^{-i}_{s},T-s)-v^{i}(X^{i}_{s-},\bar{X}^{-i}_{s-},T-s)\right],$
where the superscript $(\cdot)^{c}$ denotes the continuous part of a process. Inserting the respective dynamics (11) and (15), we get
$\begin{split}&v^{i}(X^{i}_{t},\bar{X}^{-i}_{t},T-t)\\ &=v^{i}(x^{i}_{0},\bar{x}_{0},T)-\int_{0}^{t}\partial_{T}v^{i}(X^{i}_{s-},\bar{X}^{-i}_{s-},T-s)ds\\ &\quad+\int_{0}^{t}\partial_{x}v^{i}(X^{i}_{s-},\bar{X}^{-i}_{s-},T-s)X^{i}_{s-}\left[(r^{i}+\hat{\phi}^{i}_{s}(0)(\kappa^{i}-r^{i}))ds+\sigma^{i}\hat{\phi}^{i}_{s}(0)dW^{i}_{s}+\sigma^{0i}\hat{\phi}^{i}_{s}(0)dW^{0}_{s}\right]\\ &\quad+\int_{0}^{t}\partial_{\bar{x}}v^{i}(X^{i}_{s-},\bar{X}^{-i}_{s-},T-s)\bar{X}^{-i}_{s-}\bigg{[}(\overline{\tau\pi})^{-i}ds+\frac{1}{2}[((\overline{\sigma^{0}\pi})^{-i})^{2}+(\overline{\sigma^{2}\pi^{2}})^{-i}]ds\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{n}\sum_{j\neq i}\sigma^{j}\pi^{j}(0)dW^{j}_{s}+(\overline{\sigma^{0}\pi})^{-i}dW^{0}_{s}\bigg{]}\\ &\quad+\int_{0}^{t}\partial_{x\bar{x}}v^{i}(X^{i}_{s-},\bar{X}^{-i}_{s-},T-s)X^{i}_{s-}\bar{X}^{-i}_{s-}\sigma^{0i}\hat{\phi}^{i}_{s}(0)(\overline{\sigma^{0}\pi})^{-i}ds\\ &\quad+\frac{1}{2}\int_{0}^{t}\partial_{xx}v^{i}(X^{i}_{s-},\bar{X}^{-i}_{s-},T-s)(X^{i}_{s-})^{2}\left[(\sigma^{i})^{2}+(\sigma^{0i})^{2}\right](\hat{\phi}^{i}_{s}(0))^{2}ds\\ &\quad+\frac{1}{2}\int_{0}^{t}\partial_{\bar{x}\bar{x}}v^{i}(X^{i}_{s-},\bar{X}^{-i}_{s-},T-s)(\bar{X}^{-i}_{s-})^{2}[((\overline{\sigma^{0}\pi})^{-i})^{2}+(\overline{\sigma^{2}\pi^{2}})^{-i}]ds\\ &\quad+\int_{[0,t]\times E}\bigg{[}v\left(X^{i}_{s-}(1+\hat{\phi}^{i}_{s}(\zeta^{i}(e))\eta^{i}(e)),\bar{X}^{-i}_{s-}(\overline{1+(\pi\circ\zeta)\eta})^{-i}(e),T-s\right)\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-v^{i}(X^{i}_{s-},\bar{X}^{-i}_{s-},T-s)\bigg{]}\big{(}\bar{N}(ds,de)+ds\otimes\nu(de)\big{)},\end{split}$
where, $\bar{N}$ is the compensated Poisson random measure $\bar{N}(ds,de)=N(ds,de)-ds\otimes\nu(de)$.
Following the martingale optimality principle, we look for conditions such that the value process is a supermartingale for any admissible $\tilde{\phi}^{i}$ and a martingale for some (then) optimal $\tilde{\phi}^{i,\star}$. This leads us to the following HJB equation:
(16) $\begin{split}&-\partial_{T}v^{i}(x,\bar{x},T)+\partial_{x}v^{i}(x,\bar{x},T)xr^{i}+\partial_{\bar{x}}v^{i}(x,\bar{x},T)\bar{x}[(\overline{\tau\pi})^{-i}+\tfrac{1}{2}[((\overline{\sigma^{0}\pi})^{-i})^{2}+(\overline{\sigma^{2}\pi^{2}})^{-i}]]\\ &\quad+\tfrac{1}{2}\partial_{\bar{x}\bar{x}}v^{i}(x,\bar{x},T)\bar{x}^{2}[((\overline{\sigma^{0}\pi})^{-i})^{2}+(\overline{\sigma^{2}\pi^{2}})^{-i}]\\ &\quad+\sup_{\varphi\in\bar{\Phi}^{i}(0)}\bigg{\{}\partial_{x}v^{i}(x,\bar{x},T)x\varphi(\kappa^{i}-r^{i}))+\partial_{x\bar{x}}v^{i}(x,\bar{x},T)x\bar{x}\sigma^{0i}\varphi(\overline{\sigma^{0}\pi})^{-i}\\ &\quad\quad\quad\quad\quad\quad\quad\quad+\tfrac{1}{2}\partial_{xx}v^{i}(x,\bar{x},T)x^{2}\big{[}(\sigma^{i})^{2}+(\sigma^{0i})^{2}\big{]}\varphi^{2}\\ &\quad\quad\quad\quad\quad\quad\quad\quad+\int_{\{\zeta^{i}=0\}}\bigg{[}v^{i}\big{(}x(1+\varphi\eta^{i}(e)),\bar{x}(\overline{1+(\pi\circ\zeta)\eta})^{-i}(e),T\big{)}-v^{i}(x,\bar{x},T)\bigg{]}\nu(de)\bigg{\}}\\ &\quad+\int_{\zeta^{i}(E)\setminus\{0\}}\sup_{\varphi\in\bar{\Phi}^{i}(z)}\bigg{\{}\int_{\{\zeta^{i}=z\}}\bigg{[}v^{i}\big{(}x(1+\varphi\eta^{i}(e)),\bar{x}(\overline{1+(\pi\circ\zeta)\eta})^{-i}(e),T\big{)}\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-v^{i}(x,\bar{x},T)\bigg{]}K^{i}(z,de)\bigg{\}}\mu^{i}(dz)=0\end{split}$
with boundary condition
$v^{i}(x,\bar{x},0)=u_{\alpha^{i}}(x\bar{x}^{-\theta^{i}})=\frac{(x\bar{x}^{-\theta^{i}})^{1-\alpha^{i}}}{1-\alpha^{i}}.$
To prepare the verification theorem for the HJB equation, we include the following auxiliary lemma.
Lemma 2.9.
For all $p>0$, there exist $L^{p}(\nu)$-integrable functions $f:E\to\mathbb{R}$ and $g:E\to\mathbb{R}$ such that for all $i\in\{1,...,n+1\}$ we have:
For all $z\in\zeta^{i}(E)\cup\{0\}$, for all $\varphi\in\bar{\Phi}^{i}(z)$
$(1+\varphi\eta^{i}(e))\leq f(e),$
and, for all admissible signal-driven strategies $\pi^{-i}=(\pi^{j})_{j\neq i}$,
$(\overline{1+(\pi\circ\zeta)\eta})^{-i}(e)\leq g(e).$
Proof.
For the first statement, observe that
$(1+\varphi\eta^{i}(e))\leq(1+\max_{x\in\bar{\Phi}^{i}(z)}|x|)(1+|\eta^{i}(e)|),$
where the right-handside is independent from $\varphi$ and its $p$th power is $\nu$-integrable by $(\mathtt{t}_{4})$. Similarly, for the second statment
$(\overline{1+(\pi\circ\zeta)\eta})^{-i}(e)$ $=\prod_{j\neq i}(1+\pi^{j}(\zeta^{j}(e))\eta^{j}(e))^{\frac{1}{n}}$
$\leq\prod_{j\neq i}\big{(}\max_{z\in\zeta^{j}(E)\cup\{0\}}(1+|\pi^{j}(z)|)^{\frac{1}{n}}\big{)}(1+|\eta^{j}(e)|)^{\frac{1}{n}}$
$\leq\max_{j\neq i}\max_{z\in\zeta^{j}(E)\cup\{0\}}\max_{x\in\bar{\Phi}^{i}(z)}(1+|x|)(1+|\eta^{j}(e)|)$
where the last line is independent from $\pi^{-i}$ and its $p$th power is $\nu$-integrable by $(\mathtt{t}_{4})$. ∎
With this result in place, we present the verification theorem for the best response problem and its proof:
Theorem 2.10.
Assume that our type assumption 2.4 holds and suppose investor $i$’s peers use admissible signal-driven strategies $\pi^{-i}=(\pi^{j})_{j\in\{1,...,n+1\}\setminus\{i\}}$.
Then the value function of the $i$th investor’s best response problem (10) is given by
(17) $v^{i}(x^{i}_{0},\bar{x}_{0},T)=u_{\alpha^{i}}(x^{i}_{0}(\bar{x}_{0})^{-\theta^{i}})e^{T(1-\alpha^{i})((1-\theta^{i})r^{i}+M_{\pi^{-i}}^{i})}$
where
(18) $\begin{split}M^{i}_{\pi^{-i}}&:=-\theta^{i}(\overline{\tau\pi})^{-i}+\tfrac{1}{2}(\theta^{i})^{2}(1-\alpha^{i})[((\overline{\sigma^{0}\pi})^{-i})^{2}+(\overline{\sigma^{2}\pi^{2}})^{-i}]\\ &\quad+\sup_{\varphi\in\bar{\Phi}^{i}(0)}\bigg{\{}(\kappa^{i}-r^{i})\varphi-\alpha^{i}\tfrac{1}{2}[(\sigma^{i})^{2}+(\sigma^{0i})^{2}]\varphi^{2}-\theta^{i}(1-\alpha^{i})\sigma^{0i}\varphi(\overline{\sigma^{0}\pi})^{-i}\\ &\quad\quad\quad\quad\quad\quad\quad+\int_{\{\zeta^{i}=0\}}\bigg{[}u_{i}\left(1+\varphi\eta^{i}(e),(\overline{1+(\pi\circ\zeta)\eta})^{-i}(e)\right)-u_{i}(1,1)\bigg{]}\nu(de)\bigg{\}}\\ &\quad+\int_{\zeta^{i}(E)\setminus\{0\}}\sup_{\varphi\in\bar{\Phi}^{i}(z)}\bigg{\{}\int_{\{\zeta^{i}=z\}}\bigg{[}u_{i}\left(1+\varphi\eta^{i}(e),(\overline{1+(\pi\circ\zeta)\eta})^{-i}(e)\right)-u_{i}(1,1)\bigg{]}K^{i}(z,de)\bigg{\}}\mu^{i}(dz)\end{split}$
is a finite constant. The function $v^{i}$ is indeed a smooth solution of the HJB equation (16) and the $i$th investor’s best response control is unique and it is again an admissible signal-driven strategy $\phi^{i,\star}:=\phi^{i,\star}_{\pi^{-i}}\in\mathcal{A}_{\text{sig}}^{i}$: when receiving no signal, investor $i$’s best response to $\pi^{-i}$ is holding in stock a fraction of total wealth
(19) $\begin{split}\phi^{i,\star}_{\pi^{-i}}(0)&=\operatorname*{arg\,max}_{\varphi\in\bar{\Phi}^{i}(0)}\bigg{\{}(\kappa^{i}-r^{i})\varphi-\alpha^{i}\tfrac{1}{2}[(\sigma^{i})^{2}+(\sigma^{0i})^{2}]\varphi^{2}-\theta^{i}(1-\alpha^{i})\sigma^{0i}\varphi(\overline{\sigma^{0}\pi})^{-i}\\ &\quad\quad\quad\quad\quad\quad\quad+\int_{\{\zeta^{i}=0\}}\bigg{[}u_{i}\left(1+\varphi\eta^{i}(e),(\overline{1+(\pi\circ\zeta)\eta})^{-i}(e)\right)-u_{i}(1,1)\bigg{]}\nu(de)\bigg{\}}\end{split}$
and, at moments when receiving a signal $z\neq 0$, changes this to a fraction of size
(20) $\phi^{i,\star}_{\pi^{-i}}(z)=\operatorname*{arg\,max}_{\varphi\in\bar{\Phi}^{i}(z)}\bigg{\{}\int_{\{\zeta^{i}=z\}}\bigg{[}u_{i}\left(1+\varphi\eta^{i}(e),(\overline{1+(\pi\circ\zeta)\eta})^{-i}(e)\right)-u_{i}(1,1)\bigg{]}K^{i}(z,de)\bigg{\}}.$
Proof.
(cf. [BK22, proof of Theorem 3.1]) Let us first argue that $M_{\pi^{-i}}^{i}\in\mathbb{R}$. For this, as by $(\mathtt{t}_{3})$ $\mu^{i}$ is a finite measure, we only need to show finiteness of the suprema in (18), for which we take a closer look at its integrands. Observe that for all $z\in\zeta^{i}(E)\cup\{0\}$, for all $\varphi\in\bar{\Phi}^{i}(z)$ we have
(21) $\left|u_{i}\left(1+\varphi\eta^{i}(e),(\overline{1+(\pi\circ\zeta)\eta})^{-i}(e)\right)\right|\leq\frac{(1+\varphi\eta^{i}(e))^{1-\alpha^{i}}}{|1-\alpha^{i}|}\big{(}(\overline{1+(\pi\circ\zeta)\eta})^{-i}(e)\big{)}^{-\theta^{i}(1-\alpha^{i})}.$
Thus, both for $\alpha^{i}\in(0,1)$ and $\alpha^{i}\in(1,\infty)$, the right-hand side exhibitis a denominator which is uniformly bounded away from zero by Lemma 2.5, and a numerator which is bouned from above by a $\nu$-integrable function, which does not depend on $\varphi$, by Lemma 2.9. Hence, both suprema in (18) are taken for continuous functions over compacts and thus are attained and finite.
Now, a simple calculation shows that $v^{i}(x,\bar{x},T)$ as defined in (17) is a $C^{2,2,1}((0,\infty)\times(0,\infty)\times[0,T])$-solution of (16) and meets the initial condition $v^{i}(x,\bar{x},0)=u_{\alpha^{i}}(x\bar{x}^{-\theta^{i}})$ for all $(x,\bar{x})\in(0,\infty)^{2}$. Note that $v$ satisfies $\lim_{\varepsilon\downarrow 0}v^{i}(\varepsilon,\bar{x},T-t)=v^{i}(0+,\bar{x},T-t)=u_{i}(0,\bar{x})$ for all $t\in[0,T]$, $\bar{x}\in(0,\infty)$, by our convention (7).
Let us argue that, for $\varepsilon>0$ fixed, $V^{i,\varepsilon}_{t}:=v^{i}(X^{i}_{t}+\varepsilon,\bar{X}^{-i}_{t},T-t)$, $t\in[0,T]$, is a supermartingale where $X^{i}$ solves the wealth dynamics (11) for some $\tilde{\phi}^{i}\in\mathcal{A}^{i}$. For this, we show that $V^{i,\varepsilon}$ is bounded from below, i.e. $V^{i,\varepsilon}\geq-C_{\varepsilon,\pi^{-i}}$ for some integrable $C_{\varepsilon,\pi^{-i}}>0$. Indeed, for $\alpha^{i}\in(0,1)$, the utility is bounded from below by zero. For $\alpha^{i}>1$, however, we ensure a uniform bound by showing that
(22) $\mathbb{E}[\sup_{t\in[0,T]}|\bar{X}^{-i}_{t}|^{p}]<\infty$
for all $p>0$. In fact, the mean wealth process $\bar{X}^{-i}$ is the product of independent stochastic exponentials and strictly positive and can be written in closed form as
(23) $\begin{split}\bar{X}^{-i}_{t}=\bar{x}_{0}\exp\bigg{(}(\overline{\tau\pi})^{-i}t&+\frac{1}{n}\sum_{j\neq i}\sigma^{j}\pi^{j}(0)W^{j}_{t}+(\overline{\sigma^{0}\pi})^{-i}W^{0}_{t}\\ &+\int_{[0,t]\times E}\log((\overline{1+(\pi\circ\zeta)\eta})^{-i}(e))N(ds,de)\bigg{)}\end{split}$
for $t\in[0,T]$. Let $p>0$ and let us argue for one component after another: First, note that
$\sup_{t\in[0,T]}\exp\big{(}p(\overline{\tau\pi})^{-i}t\big{)}$
does not involve any randomness and is trivially finite. Secondly, as the maximum Brownian motion has finite exponential moments, it is easy to see that
$\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\exp\bigg{(}p\cdot\frac{1}{n}\sum_{j\neq i}\sigma^{j}\pi^{j}(0)W^{j}_{t}\bigg{)}\bigg{]}<\infty\quad\text{ and }\quad\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\exp\big{(}p\cdot(\overline{\sigma^{0}\pi})^{-i}W^{0}_{t}\big{)}\bigg{]}<\infty.$
Hence, we are left to observe that
$\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\exp\bigg{\{}\int_{[0,t]\times E}$ $p\log((\overline{1+(\pi\circ\zeta)\eta})^{-i}(e))N(ds,de)\bigg{\}}\bigg{]}$
$\leq\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\int_{[0,t]\times E}((\overline{1+(\pi\circ\zeta)\eta})^{-i})(e))^{p}N(ds,de)\bigg{]}<\infty,$
but, since the integral against the Poisson random measure admits almost surely finitely many jumps on any finite time interval, we only need to check whether the jumps have finite expectation with respect to $\nu$ and this immediately follows from Lemma 2.9.
Thus, (22) holds and together with $V^{i,\varepsilon}$ having nonnegative drift, which follows from our arguments leading to the HJB equation, this implies that $V^{i,\varepsilon}$ is a supermartingale. In particular, this allows us to observe
$\mathbb{E}[u_{i}(X^{i}_{T},\bar{X}^{-i}_{T})]\leq\mathbb{E}[V^{i,\varepsilon}_{T}]\leq\mathbb{E}[V^{i,\varepsilon}_{0}]=v^{i}(x^{i}_{0}+\varepsilon,\bar{x}_{0},T).$
By sending $\varepsilon\downarrow 0$, we even have
$\mathbb{E}[u_{i}(X^{i}_{T},\bar{X}^{-i}_{T})]\leq v^{i}(x^{i}_{0},\bar{x}_{0},T).$
So the function on the right-hand side of (17) yields an upper bound for the value of investor $i$’s best response problem (10).
It remains to show that $V^{i,\star}_{t}:=v^{i}(X^{i,\star}_{t},\bar{X}^{-i}_{t},T-t)$ is a martingale for the wealth process $X^{i,\star}:=X^{\phi^{i,\star}}$ given the strategy $\phi^{i,\star}=\phi^{i,\star}_{\pi^{-i}}$. Then, this strategy is optimal since we can conclude
$v^{i}(x^{i}_{0},\bar{x}_{0},T)$ $=\mathbb{E}[u_{i}(X^{i,\star}_{T},\bar{X}^{-i}_{T})]\leq\sup_{\phi}\mathbb{E}[u_{i}(X^{i}_{T},\bar{X}^{-i}_{T})]=v^{i}(x^{i}_{0},\bar{x}_{0},T).$
We note that $\phi^{i,\star}$ is signal-driven (and admissible) in the sense of Definition 2.7 and that the use of $=$ instead of $\in$ is justified since the $\operatorname*{arg\,max}$ in both (19) and (20) are singletons. This follows immediately from the fact that the target functions are strictly concave in $\varphi$.
Further, our arguments leading to the HJB equation (16), show that $V^{i,\star}$ has zero drift:
$dV^{i,\star}_{t}=V^{i,\star}_{t-}(1-\alpha^{i})$ $\bigg{[}\sigma^{i}\phi^{i,\star}(0)dW^{i}_{t}+\sigma^{0i}\phi^{i,\star}(0)dW^{0}_{t}+\frac{1}{n}\sum_{j\neq i}\sigma^{j}\pi^{j}(0)dW^{j}_{t}+(\overline{\sigma^{0}\pi})^{-i}dW^{0}_{t}$
$\quad+\int_{E}\bigg{[}u_{i}\big{(}1+\phi^{i,\star}(\zeta^{i}(e))\eta^{i}(e),(\overline{1+(\pi\circ\zeta)\eta})^{-i}(e)\big{)}-u_{i}(1,1)\bigg{]}\bar{N}(dt,de)\bigg{]}.$
In particular, $V^{i,\star}$ is the product of independent exponentials. The stochastic exponentials with respect to the Brownian motions are martingales, since $\sigma^{i},\sigma^{0i},\sigma^{j},\sigma^{0j}$, $(1-\alpha^{i})$, $\pi^{j}(0)$ and $\phi^{i,\star}_{\pi}(0)$ are constants, and the exponential with respect to the compensated compound Poisson process is a martingale, because (cf. [Sat99, Theorem 33.2])
$\int_{E}\big{[}(1+\phi^{i,\star}(\zeta^{i}(e))\eta^{i}(e))^{\frac{1-\alpha^{i}}{2}}((\overline{1+(\pi\circ\zeta)\eta})^{-i}(e))^{\frac{-\theta^{i}(1-\alpha^{i})}{2n}}-1\big{]}^{2}\nu(de)$
$\quad\quad\leq\int_{E}\big{[}(1+\phi^{i,\star}(\zeta^{i}(e))\eta^{i}(e))^{1-\alpha^{i}}((\overline{1+(\pi\circ\zeta)\eta})^{-i}(e))^{\frac{-\theta^{i}(1-\alpha^{i})}{n}}+1\big{]}\nu(de)$
is ensured to be finite by $M_{\pi^{-i}}^{i}$ being finite. Thus, $V^{i,\star}$ is a martingale and the proof is complete. ∎
2.4. Existence of a Nash Equilibrium
Having found the best response for a single investor interacting with peers employing signal-driven strategies, we are now in a position to demonstrate the existence of an equilibrium.
Theorem 2.11.
Under Assumption 2.4, there exists a signal-driven Nash equilibrium $(\pi^{1,\star},....,\pi^{n+1,\star})\in\mathcal{A}_{\text{sig}}^{1}\times...\times\mathcal{A}_{\text{sig}}^{n+1}$.
The proof is based on Schauder’s Fixed Point Theorem which requires a continuous self-mapping of a compact and convex domain. For our best response map above (19) and (20), these requirements are difficult to meet in general, but our Assumption $(\mathtt{t}_{3})$ makes this tractable: There are only finitely many distinct signals to account for and so the space of admissible signal-driven strategies $(\pi^{1},....,\pi^{n+1})\in\mathcal{A}_{\text{sig}}^{1}\times...\times\mathcal{A}_{\text{sig}}^{n+1}$ can be identified with a vector
$(\pi^{1},...,\pi^{n+1})\in\bigtimes_{i=1}^{n+1}\bigtimes_{z\in\zeta^{i}(E)\cup\{0\}}\bar{\Phi}^{i}(z)=:\bar{\Phi}.$
The set $\bar{\Phi}$ is compact and convex by construction and thus serves as the domain and image set of the desired fixed point mapping.
Proof.
A sufficient condition for a candidate $\pi=(\pi^{1},...,\pi^{n+1})\in\mathcal{A}_{\text{sig}}^{1}\times...\times\mathcal{A}_{\text{sig}}^{n+1}$ to be a signal-driven Nash equilibrium is to satisfy $\pi^{i}=\phi^{i,\star}_{\pi^{-i}}\text{ for all }i\in\{1,...,n+1\}$, where $\phi^{i,\star}_{\pi^{-i}}$ is described in Theorem 2.10. Thus, the equilibrium is characterized by a fixed point of
$\gamma:\bar{\Phi}\to\bar{\Phi},\quad\pi=(\pi^{1},...,\pi^{n+1})\mapsto\gamma(\pi):=(\gamma(\pi)^{1},...,\gamma(\pi)^{n+1}):=(\phi^{1,\star}_{\pi^{-1}},...,\phi^{n+1,\star}_{\pi^{-(n+1)}}).$
Let us argue for continuity of $\gamma$. By Lemma 2.9, we can use dominated convergence to find that the target functions in both (19) and (20) are continuous in $(\varphi,\pi)$. Thus, by Berge’s Maximum Theorem (cf. [AB06, Theorem 17.31]), $\phi^{i,\star}_{\pi^{-i}}$ is upper-hemicontinuous in $\pi$. In fact, since by strict concavity of our utility functions the $\operatorname*{arg\,max}$ operators in (19) and (20) always yield singletons, the maximizers are actually continuous functions. As a consequence, $\gamma$ is indeed continuous in $\pi$.
Finally, since $\bar{\Phi}$ is non-empty, compact and convex and $\gamma:\bar{\Phi}\to\bar{\Phi}$ is continuous, we can use Schauder’s Fixed Point Theorem, cf. [Sch30], to conclude that there exists $\pi^{\star}\in\bar{\Phi}$ with $\pi^{\star}=\gamma(\pi^{\star})$. ∎
In summary, we have demonstrated the existence of a signal-driven Nash equilibrium, which also qualifies as a Nash equilibrium within the broader class of admissible and not necessarily signal-driven strategies.
2.5. Common vs. Idiosyncratic Noise in Aggregate Wealth
An important question arises: how can this framework transition to a corresponding mean field game? This transition is particularly motivated by the fact that, in such a mean field setting a single individual’s actions do not impact the environment of others and, so, the signal-driven strategies that we are focussing on are an even more natural concept as the average wealth at the core of the interactions between agents cannot be influenced anymore by every single investor. One way to interpret the mean field setting is to consider it as the limit of the multi-agent game for the number of agents $n\to\infty$.
Before moving on to a rigorous formulation of the mean field game in Section 3 below, we provide an appetizer discussion and deliver a result on the limit of the average wealth for future reference.
Let us assume that the types of our investors $j=1,2,\dots$ are obtained from a corresponding sequence $\mathtt{t}^{1}$, $\mathtt{t}^{2}$,… in $\mathtt{T}$ generated by independently and identically distributed samples from a probability $\mathbb{T}=\mathbb{T}(d\mathtt{t})$ which describes the distribution of types among the population of investors. We denote a generic sample from $\mathbb{T}$ by
$\mathtt{t}=(x_{0}^{\mathtt{t}},r^{\mathtt{t}},\kappa^{\mathtt{t}},\sigma^{\mathtt{t}},\sigma^{0\mathtt{t}},\eta^{\mathtt{t}},\zeta^{\mathtt{t}},\alpha^{\mathtt{t}},\theta^{\mathtt{t}})\in\mathtt{T}$
and with some abuse of notation consider
$\mathtt{t}^{j}=(x_{0}^{j},r^{j},\kappa^{j},\sigma^{j},\sigma^{0j},\eta^{j},\zeta^{j},\alpha^{j},\theta^{j})=(x_{0}^{\mathtt{t}^{j}},r^{\mathtt{t}^{j}},\kappa^{\mathtt{t}^{j}},\sigma^{\mathtt{t}^{j}},\sigma^{0\mathtt{t}^{j}},\eta^{\mathtt{t}^{j}},\zeta^{\mathtt{t}^{j}},\alpha^{\mathtt{t}^{j}},\theta^{\mathtt{t}^{j}})\in\mathtt{T}\quad\text{ for }j\in\mathbb{N}.$
Each investor $j=1,2,\dots$ is equipped with her own independent idiosyncratic Brownian motion $W^{j}$; the Brownian motion $W^{0}$ specifies common Brownian noise shared by all agents.
The specification of common and idiosyncratic jump noise is less obvious. A particular challenge in this context is the specification of a framework which allows for idiosyncratic signals on (partially) common price shocks, a key feature of our model, which, to the best of our knowledge, has not been addressed in the literature so far. Indeed, while jump times of the Poisson random measure $N(dt,de)$ should be common to all investors, it is important to address the question how the marks set by $N$ represent common and idiosyncratic noise. For instance, we want to allow for the possibility that each of our investors receives a signal on common impending price shocks which is affected by idiosyncratic noise. For instance, an investor may have a type $\mathtt{t}\in\mathtt{T}$ that affords her warnings about impending price shocks with some probability $p^{\mathtt{t}}$. While the price shock may be common to all investors and the warning probability $p^{\mathtt{t}}$ will be shared by her fellow investors of type $\mathtt{t}$, the actual heads-or-tails outcome of her idiosyncratic coin flip experiment for the shock warning may be independent from that of everyone else. The distinction between common and idiosyncratic noise becomes particularly crucial in the mean field framework. Here, idionsyncratic noise can be averaged out by a law of large numbers to yield a deterministic quantity while common noise persists as a random quantity affecting everyone.
To allow for a rigorous formalization of such settings with idosyncratic and common jump noise, let us consider the mark space as an infinite product space
$E:=E^{\mathtt{c}}\times E^{\otimes\mathtt{i}}\text{ for }E^{\otimes\mathtt{i}}:=\bigtimes_{j\in\mathbb{N}}E^{\mathtt{i},j}$
for Polish spaces $E^{\mathtt{c}}$ and identical $E^{\mathtt{i},1}=E^{\mathtt{i},2}=\dots=:E^{\mathtt{i}}$ with generic elements
$e=(e^{\mathtt{c}},e^{\mathtt{i}})=(e^{\mathtt{c}},e^{\mathtt{i},1},e^{\mathtt{i},2},...)\in E.$
The independent Poisson random measure $N=N(dt,de)=N(dt,de^{\mathtt{c}},de^{\mathtt{i}})$ has intensity measure
$\nu:=\lambda(\nu^{\mathtt{c}}\otimes\nu^{\otimes\mathtt{i}}),$
where $\nu^{\mathtt{c}}$ is a probability measure on $E^{\mathtt{c}}$ and
$\nu^{\otimes\mathtt{i}}(de^{\mathtt{i}}):=\bigotimes_{j\in\mathbb{N}}\nu^{\mathtt{i}}(de^{\mathtt{i},j})$
generates independently for each investor identically distributed idiosyncratic noise marks, each with law $\nu^{\mathtt{i}}$ on $E^{\mathtt{i}}$. In particular, this leads to jumps and signals with both common and idiosyncratic components for each investor $j=1,2,\dots$ via $\eta^{j}(e^{\mathtt{c}},e^{\mathtt{i}})=\eta^{\mathtt{t}^{j}}(e^{\mathtt{c}},e^{\mathtt{i}})=\eta^{\mathtt{t}^{j}}(e^{\mathtt{c}},e^{\mathtt{i},j})$ and $\zeta^{j}(e^{\mathtt{c}},e^{\mathtt{i}})=\zeta^{\mathtt{t}^{j}}(e^{\mathtt{c}},e^{\mathtt{i}})=\zeta^{j}(e^{\mathtt{c}},e^{\mathtt{i},j})$ where $\eta^{j}$ and $\zeta^{j}$ are the jump- and signal maps on $E^{\mathtt{c}}\times E^{\mathtt{i}}$ for an investor of type $\mathtt{t}^{j}\in\mathtt{T}$.
Of course, the actual signals $Z_{t_{k}}^{j}=\zeta^{\mathtt{t}^{j}}(e^{\mathtt{c}}_{k},e^{\mathtt{i},j}_{k})$ that each individual investor $j=1,2,\dots$ may, at the various mark times $t_{k}$, receive are exposed to both the respective common mark $e^{\mathtt{c}}_{k}$ and the investor specific mark $e^{\mathtt{i},j}_{k}$. As a result, even investors of the same type may make differing investment decisions and figuring out the limiting geometric average wealth when their number $n$ goes to infinity has to account for this:
Lemma 2.12.
Suppose each investor $j=1,2,\dots$ follows a signal-based strategy $\pi^{j}=\pi^{\mathtt{t}^{j}}\in\mathcal{A}^{j}_{\text{sig}}$ determined solely by her respective type $\mathtt{t}^{j}$, determined as i.i.d. sample from the type distibution $\mathbb{T}(d\mathtt{t})$. Moreover, suppose the following integrability assumptions hold:
$\begin{split}&\int_{\mathtt{T}}|\log(x_{0}^{\mathtt{t}})|\mathbb{T}(d\mathtt{t})<\infty,\\ &\int_{\mathtt{T}}\big{|}r^{\mathtt{t}}+\pi^{\mathtt{t}}(0)(\kappa^{\mathtt{t}}-r^{\mathtt{t}})-\frac{1}{2}((\sigma^{\mathtt{t}})^{2}+(\sigma^{0\mathtt{t}})^{2})(\pi^{\mathtt{t}}(0))^{2}\big{|}\mathbb{T}(d\mathtt{t})<\infty,\\ &\int_{\mathtt{T}}|\sigma^{0\mathtt{t}}\pi^{\mathtt{t}}(0)|\mathbb{T}(d\mathtt{t})<\infty\end{split}$
and for $e^{\mathtt{c}}\in E^{\mathtt{c}}$
$\int_{\mathtt{T}\times E^{\mathtt{i}}}\log\big{(}1+\pi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}})\big{)}\nu^{i}(de^{\mathtt{i}})\otimes\mathbb{T}(d\mathtt{t})<\infty.$
Then, as the number of investors $n$ converges to $\infty$, the average wealth $\bar{X}^{-i}_{t}$ of each investor’s $i\in\mathbb{N}$ peers converges for all $t\in[0,T]$ almost surely to
$\lim_{n\to\infty}\bar{X}^{-i}_{t}=\bar{x}_{0}\exp$$\bigg{\{}\overline{\mathtt{t}\pi}\ t+\overline{\sigma^{0}\pi}\ W^{0}_{t}+\int_{[0,t]\times E^{\mathtt{c}}}\log(\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}}))N(ds,de^{\mathtt{c}},E^{\mathtt{i}})\bigg{\}}$
for
(24) $\begin{split}\bar{x}_{0}&:=\exp\int_{\mathtt{T}}\log(x_{0}^{\mathtt{t}})\mathbb{T}(d\mathtt{t})\\ \overline{\tau\pi}&:=\int_{\mathtt{T}}\left(r^{\mathtt{t}}+\pi^{\mathtt{t}}(0)(\kappa^{\mathtt{t}}-r^{\mathtt{t}})-\frac{1}{2}((\sigma^{\mathtt{t}})^{2}+(\sigma^{0\mathtt{t}})^{2})(\pi^{\mathtt{t}}(0))^{2}\right)\mathbb{T}(d\mathtt{t})\\ \overline{\sigma^{0}\pi}&:=\int_{\mathtt{T}}\sigma^{0\mathtt{t}}\pi^{\mathtt{t}}(0)\mathbb{T}(d\mathtt{t})\end{split}$
and
(25) $\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}}):=\exp\left\{\int_{\mathtt{T}\times E^{\mathtt{i}}}\log\big{(}1+\pi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}})\big{)}\nu^{i}(de^{\mathtt{i}})\otimes\mathbb{T}(d\mathtt{t})\right\},\quad e^{\mathtt{c}}\in E^{\mathtt{c}}.$
Proof.
For $n\in\mathbb{N}$, $i=1,\dots,n$, and $t\in[0,T]$ consider $\bar{X}^{-i}$ from (14) for the specified noise structure:
$\bar{X}^{-i}_{t}=\bigg{(}\prod_{j\neq i}x_{0}^{j}\bigg{)}^{\frac{1}{n}}\exp\bigg{\{}t\frac{1}{n}\sum_{j\neq i}\big{[}r^{j}$ $+\pi^{j}(0)(\kappa^{j}-r^{j})-\frac{1}{2}((\sigma^{j})^{2}+(\sigma^{0j})^{2})(\pi^{j}(0))^{2}\big{]}+\frac{1}{n}\sum_{j\neq i}\sigma^{j}\pi^{j}(0)W^{j}_{t}$
$+\frac{1}{n}\sum_{j\neq i}\sigma^{0j}\pi^{j}(0)W^{0}_{t}+\sum_{t_{k}\leq t}\frac{1}{n}\sum_{j\neq i}\log(1+\pi^{j}(\zeta^{j}(e^{\mathtt{c}}_{k},e^{\mathtt{i},j}_{k}))\eta^{j}(e^{\mathtt{c}}_{k},e^{\mathtt{i},j}_{k}))\bigg{\}}.$
By the law of large numbers, we immediately find
$\lim_{n\to\infty}\bigg{(}\prod_{j\neq i}x_{0}^{j}\bigg{)}^{\frac{1}{n}}=\lim_{n\to\infty}\exp\bigg{(}\frac{1}{n}\sum_{j\neq i}\log(x^{j}_{0})\bigg{)}=\exp\int_{\mathtt{T}}\log(x_{0}^{\mathtt{t}})\mathbb{T}(d\mathtt{t}).$
Similarly, we find convergence of
$\lim_{n\to\infty}\frac{1}{n}\sum_{j\neq i}\big{[}r^{j}$ $+\pi^{j}(0)(\kappa^{j}-r^{j})-\frac{1}{2}((\sigma^{j})^{2}+(\sigma^{0j})^{2})(\pi^{j}(0))^{2}\big{]}$
$=\int_{\mathtt{T}}\left(r^{\mathtt{t}}+\pi^{\mathtt{t}}(0)(\kappa^{\mathtt{t}}-r^{\mathtt{t}})-\frac{1}{2}((\sigma^{\mathtt{t}})^{2}+(\sigma^{0\mathtt{t}})^{2})(\pi^{\mathtt{t}}(0))^{2}\right)\mathbb{T}(d\mathtt{t}).$
For the term aggregating the independent idiosyncratic Brownian motions $W^{j}$, we find
$\lim_{n\to\infty}\frac{1}{n}\sum_{j\neq i}\sigma^{j}\pi^{j}(0)W^{j}_{t}=0.$
For the term involving the common noise Brownian motion $W^{0}$, we get the average
$\lim_{n\to\infty}\frac{1}{n}\sum_{j\neq i}\sigma^{0j}\pi^{j}(0)W^{0}_{t}=W^{0}_{t}\int_{\mathtt{T}}\sigma^{0\mathtt{t}}\pi^{\mathtt{t}}(0)\mathbb{T}(d\mathtt{t}).$
Lastly, for the jump term, we note that all investors have in common the mark times $t_{k}$ and the noise marks $e^{\mathtt{c}}_{k}$, while the $e^{\mathtt{i},j}_{k}$ are i.i.d. for $k,j\in\mathbb{N}$ with distribution $\nu^{\mathtt{i}}$. The law of large numbers thus yields the almost sure convergence
$\lim_{n\to\infty}\sum_{t_{k}\leq t}\frac{1}{n}\sum_{j\neq i}$ $\log(1+\pi^{j}(\zeta^{j}(e^{\mathtt{c}}_{k},e^{\mathtt{i},j}_{k}))\eta^{j}(e^{\mathtt{c}}_{k},e^{\mathtt{i},j}_{k}))$
$=\sum_{t_{k}\leq t}\int_{E^{\mathtt{i}}\times\mathtt{T}}\log\big{(}1+\pi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e^{\mathtt{c}}_{k},e^{\mathtt{i}}))\eta^{\mathtt{t}}(e^{\mathtt{c}}_{k},e^{\mathtt{i}})\big{)}\nu^{\mathtt{i}}(de^{\mathtt{i}})\otimes\mathbb{T}(d\mathtt{t})$
$=\int_{[0,t]\times E^{\mathtt{c}}}\bigg{(}\int_{E^{\mathtt{i}}\times\mathtt{T}}\log\big{(}1+\pi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}})\big{)}\nu^{\mathtt{i}}(de^{\mathtt{i}})\otimes\mathbb{T}(d\mathtt{t})\bigg{)}N(ds,de^{\mathtt{c}},E^{\mathtt{i}}).$
Putting everything together, we find the desired result. ∎
It is an intriguing question whether, when adding more and more investors in an i.i.d. fashion, the correspodning Nash equilbria also converge to a mean field equilibrium, whose definition we provide and existence we establish in the following Section 3. Moreover, one could ask whether the average wealth of investors employing equilibrium strategies in a multi-agent game converges to the average wealth in the mean field game under its equilibrium strategy. We have to leave these questions for future research. The main reason is that we cannot even prove uniqueness of equilibria; see Remark 3.14 where we discuss the pertaining difficulties.
3. The Mean Field Game
In this section we give a mean field formulation for interactive utility maximization with jump signals. Section 3.1 provides a model for the representative investor which leads to a canonical notion of mean field equilibrium in signal-driven strategies. Section 3.2 provides the best response function which maps a proposed signal-driven stragegy used in the mean field environment to the representative investor’s best response. Section 3.3 then proves existence of a mean field equilibrium, preparing the ground for the numerical case studies carried out in the next Section 4.
3.1. The Representative Investor
Guided by the considerations of the average wealth in Section 2.5, let us assume that our representative investor’s type is determined by a probability $\mathbb{T}=\mathbb{T}(d\mathtt{t})$ on the type space
$\mathtt{T}=(0,\infty)\times\mathbb{R}^{2}\times[0,\infty)\times[0,\infty)\times L^{0}(E)\times L^{0}(E)\times((0,\infty)\setminus\{1\})\times[0,1]$
with its Borel $\sigma$-field $\mathcal{T}:=\mathcal{B}(\mathtt{T})$. As this type choice should be independent from the other sources of randomness in our system, we pass from our original probability space $(\Omega,\mathfrak{A},\mathbb{P})$ supporting independent Brownian motions $W$, $W^{0}$ along with an independent marked Poisson point process $N$ to
$\bar{\Omega}:=\mathtt{T}\times\Omega,\;\bar{\mathfrak{A}}:=\mathcal{T}\otimes\mathfrak{A}\text{ and }\bar{\mathbb{P}}:=\mathbb{T}\otimes\mathbb{P}.$
The processes $W$, $W^{0}$, $N$ are lifted to $\bar{\Omega}$ via the canonical projection $\bar{\Omega}\to\Omega$.
Similarly, we consider the projection $\tau(\bar{\omega})=\tau(\mathtt{t},\omega)=\mathtt{t}\in\mathtt{T}$ defined for a generic $\bar{\omega}=(\mathtt{t},\omega)\in\mathtt{T}\times\Omega=\bar{\Omega}$. The components of this projection $\tau$ implicitly define the population’s charateristics
$\tau=(x_{0},r,\kappa,\sigma,\sigma^{0},\eta,\zeta,\alpha,\theta)$
as measurable mappings on $\bar{\Omega}$ To emphasize their dependence on the type, we sometimes will denote $x_{0}=x_{0}^{\tau}$, i.e.  $x_{0}^{\tau}(\bar{\omega})=x_{0}^{\mathtt{t}}$, where the latter is the initial wealth compoment stored in the type vector $\mathtt{t}\in\mathtt{T}$; we proceed similarly for $r=r^{\tau}$, $\kappa=\kappa^{\tau}$, etc.
Idiosyncratic and common jump noise can be introduced in analogy to our setting in Section 2.5 except that we only need idiosyncratic and common noise for one representative investor. We thus consider a mark space in product form
$E:=E^{\mathtt{c}}\times E^{\mathtt{i}}$
with two Polish spaces $E^{\mathtt{c}}$, $E^{\mathtt{i}}$, equipped with their respective Borel-$\sigma$-fields $\mathcal{E}^{\mathtt{c}}=\mathcal{B}(E^{\mathtt{c}})$ and $\mathcal{E}^{i}=\mathcal{B}(E^{\mathtt{i}})$. We write
$\nu:=\lambda(\nu^{\mathtt{c}}\otimes\nu^{i}),$
as the product of two probability measures $\nu^{\mathtt{c}}$ and $\nu^{i}$ scaled with the jump rate $\lambda\in(0,\infty)$. Elements of $E$ will be generically denoted by $e=(e^{\mathtt{c}},e^{\mathtt{i}})\in E^{\mathtt{c}}\times E^{\mathtt{i}}=E$ and noise such as price shocks that commonly affect all investors will only depend on $e$ via its $e^{\mathtt{c}}$ component; idiosyncratic noise resulting from a mark $e$ such as shock warnings will also depend on $e^{\mathtt{i}}$, a source of randomness that we will use to allow for individual noise in signals that should be thought of as independent between agents.
Let us also lift the information flows generated by the drivers. That is, denoting by $\mathcal{N}$ the set of $\bar{\mathbb{P}}$-negligible events, we have
¯$\bar{\mathcal{F}}^{W,W^{0},N}_{t}:=\sigma((W_{s},W^{0}_{s},N([0,s]\times A\times B)):s\in[0,t],A\in\mathcal{E}^{\mathtt{c}},B\in\mathcal{E}^{i})\vee\mathcal{N}.$
Supplemented by the information generated by the type $\tau$ this gives us the full information filtration
¯$\bar{\mathcal{F}}_{t}:=\bar{\mathcal{F}}^{W,W^{0},N}_{t}\vee\sigma(\tau),\quad t\in[0,T].$
Also, denote by $\mathcal{F}^{\text{0}}=(\mathcal{F}^{0}_{t})_{t\in[0,T]}$ the common noise filtration
$\mathcal{F}^{0}_{t}:=\sigma((W^{0}_{s},N([0,s]\times A\times E^{\mathtt{i}})):s\in[0,t],A\in\mathcal{E}^{\mathtt{c}})\vee\mathcal{N}.$
An immediate, albeit fundamental consequence is that any ¯$\bar{\mathcal{F}}_{0}$-measurable random variable is actually $\sigma(\tau)$-measurable and independent of $N$, $W^{0}$ and $W$ under $\bar{\mathbb{P}}$. Additionally,
$\eta^{\tau}(\bar{\omega})(e)=\eta^{\mathtt{t}}(e)\text{ and }\zeta^{\tau}(\bar{\omega})(e)=\zeta^{\mathtt{t}}(e)$
are jointly measurable in their arguments.
As in the multi-agent game, investors will obtain information from their signal process. In the present mean field setting, we define the representative investor’s signal process as
$Z_{t}=\int_{\{t\}\times E}\zeta(e)N(ds,de)\quad\text{ for }t\in[0,T].$
Observe that this signal process is type-dependent as
$Z(\bar{\omega})=Z^{\tau(\bar{\omega})}(\bar{\omega})=Z^{\mathtt{t}}(\omega):=\left(\int_{\{t\}\times E}\zeta^{\mathtt{t}}(e)N(\omega)(ds,de)\right)_{t\in[0,T]}\text{ for }\bar{\omega}=(\mathtt{t},\omega)\in\mathtt{T}\times\Omega=\bar{\Omega}.$
In the following, we rely on the type-dependent disintegration of the jump intensity measure: For all $\mathtt{t}\in\mathtt{T}$, we have
(26) $\nu(de\cap\{\zeta^{\mathtt{t}}\neq 0\})=\int_{\mathbb{R}\setminus\{0\}}K^{\mathtt{t}}(z,de)\mu^{\mathtt{t}}(dz),\quad\mu^{\mathtt{t}}:=\nu\circ(\zeta^{\mathtt{t}})^{-1},$
where $\mathtt{t}\mapsto\mu^{\mathtt{t}}(\cdot)$ and $\mathtt{t}\mapsto K^{\mathtt{t}}(z,\cdot)$ are $\mathcal{T}$-measurable. This allows us to introduce
$\underline{\eta}^{\mathtt{t}}(z):=K^{\mathtt{t}}(z,\cdot)\text{-ess inf }\eta^{\mathtt{t}},$
$\bar{\eta}^{\mathtt{t}}(z):=K^{\mathtt{t}}(z,\cdot)\text{-ess sup}\eta^{\mathtt{t}},$
$\underline{\eta}^{\mathtt{t}}(0):=\nu(\cdot\cap\{\zeta^{\mathtt{t}}=0,\eta^{\mathtt{t}}\neq 0\})\text{-ess inf }\eta^{\mathtt{t}}\text{ and }$
$\bar{\eta}^{\mathtt{t}}(0):=\nu(\cdot\cap\{\zeta^{\mathtt{t}}=0,\eta^{\mathtt{t}}\neq 0\})\text{-ess sup }\eta^{\mathtt{t}}$
for all $\mathtt{t}\in\mathtt{T}$ and $z\in\mathbb{R}\setminus\{0\}$.
By the same reasoning as in the multi-agent game, we obtain both the no-arbitrage condition
(27) $\underline{\eta}^{\mathtt{t}}(z)<0<\bar{\eta}^{\mathtt{t}}(z)\text{ for }z\in\zeta^{\mathtt{t}}(E)\setminus\{0\},$
and the admissible positions ensuring non-negative wealth: In case $\nu(\zeta^{\mathtt{t}}=0,\eta^{\mathtt{t}}\neq 0)>0$ when unsignaled jumps are possible, the default positions without signal can be chosen from
$\Phi^{\mathtt{t}}(0):=\begin{cases}[-1/\bar{\eta}^{\mathtt{t}}(0),-1/\underline{\eta}^{\mathtt{t}}(0)],\quad&\text{ if }\underline{\eta}^{\mathtt{t}}(0)<0<\bar{\eta}^{\mathtt{t}}(0),\\ [-1/\bar{\eta}^{\mathtt{t}}(0),+\infty),\quad&\text{ if }0\leq\underline{\eta}^{\mathtt{t}}(0)\leq\bar{\eta}^{\mathtt{t}}(0),\\ (-\infty,-1/\underline{\eta}^{\mathtt{t}}(0)],\quad&\text{ if }\underline{\eta}^{\mathtt{t}}(0)\leq\bar{\eta}^{\mathtt{t}}(0)\leq 0;\end{cases}$
if $\nu(\zeta^{\mathtt{t}}=0,\eta^{\mathtt{t}}\neq 0)=0$, then all jumps are signaled and so $\Phi^{\mathtt{t}}(0):=\mathbb{R}$. For a non-zero signal we have
$\Phi^{\mathtt{t}}(z):=[-1/\bar{\eta}^{\mathtt{t}}(z),-1/\underline{\eta}^{\mathtt{t}}(z)]\text{ for }z\in\mathbb{R}\setminus\{0\}.$
We make the following standing assumption on the distribution of investor types.
Assumption 3.1.
The type distribution $\mathbb{T}$ is such that
  1. $(\mathtt{T}_{1})$
    the stock prices are nonnegative, i.e. $\nu(\{\eta^{\mathtt{t}}<-1\})=0$ for $\mathbb{T}$-almost every $\mathtt{t}\in\mathtt{T}$,
  2. $(\mathtt{T}_{2})$
    the no-arbitrage condition, i.e. $\underline{\eta}^{\mathtt{t}}(z)<0<\bar{\eta}^{\mathtt{t}}(z)\text{ for $\mu^{\mathtt{t}}$-a.e. }z\in\zeta^{\mathtt{t}}(E)\setminus\{0\}$, holds for $\mathbb{T}$-almost every $\mathtt{t}\in\mathtt{T}$,
  3. $(\mathtt{T}_{3})$
    each investor type can receive only finitely many different signals, i.e. $|\zeta^{\mathtt{t}}(E)|<\infty$ for $\mathbb{T}$-almost every $\mathtt{t}\in\mathtt{T}$, and
  4. $(\mathtt{T}_{4})$
    the market parameters and the risk aversion are bounded, i.e.
    $\mathbb{T}(d\mathtt{t})\text{-}\operatorname*{ess\,sup}(|r^{\mathtt{t}}|+|\kappa^{\mathtt{t}}|+\sigma^{\mathtt{t}}+\sigma^{0\mathtt{t}}+\alpha^{\mathtt{t}})<\infty$
    and the initial wealth allows for a finite geometric average
    $\log x^{\mathtt{t}}_{0}\in L^{1}(\mathbb{T}(d\mathtt{t})).$
  5. ($\mathtt{T}_{5}$)
    jump bounds are uniformly bounded away from zero: There exists $c_{0}>0$ such that for $\mathbb{T}$-almost every $\mathtt{t}\in\mathtt{T}$
    $\bar{\eta}^{\mathtt{t}}(z)\geq c_{0}>0\text{ and }\underline{\eta}^{\mathtt{t}}(z)\leq-c_{0}<0\quad\text{ for all }z\in\mathbb{R}$
  6. $(\mathtt{T}_{6})$
    if only finitely many distinct types can occur (i.e. if $|\operatorname*{supp}\mathbb{T}|<\infty$), jump sizes have finite moments in the sense that
    $\int_{E}(1+\eta^{\mathtt{t}}(e))^{p}\nu(de)<\infty\text{ for all }p>0,\ \mathtt{t}\in\operatorname*{supp}\mathbb{T};$
    otherwise, we need uniform boundedness in the sense that
    $\mathbb{T}(d\mathtt{t})\otimes\nu(de)\text{-}\operatorname*{ess\,sup}\eta^{\mathtt{t}}(e)<\infty.$
Let us review these assumptions. First of all, $(\mathtt{T}_{1})$-$(\mathtt{T}_{3})$ are consistent with our assumptions in the multi-agent game, see Assumption 2.4. Further, $(\mathtt{T}_{4})$ will give us bounds on the geometric mean field wealth by ensuring that market parameters do not differ too much between different types of investors. It will hold immediately if $\mathbb{T}$ distinguishes only finitely many types. Similarly, ($\mathtt{T}_{5}$) will ensure ‘nice’ properties of the mean field jumps, as it yields uniform bounds for admissible positions of all investor types, see (i) in Lemma 3.8 below. This will be particularly important in the construction of a suitable fixed point mapping in our proof for existence of mean field equilibria, see Lemma 3.12. Lastly, $(\mathtt{T}_{6})$ ensures sufficient integrability of stock price jumps, a condition also needed in the multi-agent game.
Let us next define the set of admissible strategies for the representative investor. For this, we will again consider compact intervals $\bar{\Phi}^{\mathtt{t}}(z)$ included in $\{0\}\cup\operatorname*{int}\Phi^{\mathtt{t}}(z)$ for all $z\in\mathbb{R}$, $\mathtt{t}\in\mathtt{T}$. However, to ensure measurability and uniformly nonnegative wealth, we specify the intervals to be of the following form: If $\mathbb{T}$ is finitely supported, we let
$\bar{\Phi}^{\mathtt{t}}(z):=\left[-\frac{1-a^{\mathtt{t}}}{\overline{\eta}^{\mathtt{t}}(z)},-\frac{1-b^{\mathtt{t}}}{\underline{\eta}^{\mathtt{t}}(z)}\right]\text{ for $z\in\mathbb{R}$, $\mathtt{t}\in\operatorname*{supp}\mathbb{T}$,}$
for $\mathcal{T}$-measurable $\mathtt{t}\mapsto a^{\mathtt{t}}$, $\mathtt{t}\mapsto b^{\mathtt{t}}\in(0,1]$. In case of infinitely many possible types, we consider
(28) $\bar{\Phi}^{\mathtt{t}}(z):=\left[-\frac{1-a}{\overline{\eta}^{\mathtt{t}}(z)},-\frac{1-b}{\underline{\eta}^{\mathtt{t}}(z)}\right]$
for all $\mathtt{t}\in\mathtt{T}$ and for type-independent $a,b\in(0,1]$.
To prepare the application of the Measurable Maximum Theorem (cf. [AB06, Theorem 18.19]) let us record the following lemma:
Lemma 3.2.
The map $\mathtt{t}\mapsto\bar{\Phi}^{\mathtt{t}}(z)$ is weakly $\mathcal{T}$-measurable in the sense of [AB06, Def. 18.1], that is, for all open subsets $B$ of $\mathbb{R}$, we have
$\{\mathtt{t}\in\mathtt{T}:\bar{\Phi}^{\mathtt{t}}(z)\cap B\neq\emptyset\}\in\mathcal{T}\text{ for any }z\in\mathbb{R}.$
Proof.
It suffices to show that for all $x,y\in\mathbb{R}$ we have
$\{\mathtt{t}\in\mathtt{T}:\bar{\Phi}^{\mathtt{t}}(z)\cap(x,y)\neq\emptyset\}\in\mathcal{T}$
for $z\in\mathbb{R}$. Thus, we get the desired measurability for both maps as soon as we have shown measurability of the interval bounds, that is, we need to show that each of the maps
$\mathtt{t}\mapsto\underline{\eta}^{\mathtt{t}}(0)=\nu(\cdot\cap\{\zeta^{\mathtt{t}}=0,\eta^{\mathtt{t}}\neq 0\})\text{-ess inf }\eta^{\mathtt{t}},$
$\mathtt{t}\mapsto\overline{\eta}^{\mathtt{t}}(0)=\nu(\cdot\cap\{\zeta^{\mathtt{t}}=0,\eta^{\mathtt{t}}\neq 0\})\text{-ess sup }\eta^{\mathtt{t}},$
$\mathtt{t}\mapsto\underline{\eta}^{\mathtt{t}}(z)=K^{\mathtt{t}}(z,\cdot)\text{-ess inf }\eta^{\mathtt{t}},$
$\mathtt{t}\mapsto\overline{\eta}^{\mathtt{t}}(z)=K^{\mathtt{t}}(z,\cdot)\text{-ess sup }\eta^{\mathtt{t}}$
is $\mathcal{T}$-measurable.
The argument for the other maps being similar, let us show that $\mathtt{t}\mapsto\overline{\eta}^{\mathtt{t}}(0)$ is $\mathcal{T}$-measurable. Noting that by Fubini-Tonelli
$\mathtt{t}\mapsto\|\eta_{+}^{\mathtt{t}}\|_{L^{p}(\nu(\cdot\cap\{\zeta^{\mathtt{t}}=0,\eta^{\mathtt{t}}\neq 0\}))}=\int_{E}\mathds{1}_{\{\zeta^{\mathtt{t}}(e)=0,\eta^{\mathtt{t}}(e)\neq 0\}}|\eta_{+}^{\mathtt{t}}(e)|^{p}\nu(de)$
is $\mathcal{T}$-measurable since the integrand is jointly $\mathcal{T}\otimes\mathcal{E}$-measurable and positive, we can write $\overline{\eta}^{\mathtt{t}}(0)$ as the pointwise limit of $\mathcal{T}$-measurable functions:
$\overline{\eta}^{\mathtt{t}}(0)=\nu(\cdot\cap\{\zeta^{\mathtt{t}}=0,\eta^{\mathtt{t}}\neq 0\})\text{-}\operatorname*{ess\,sup}\eta^{\mathtt{t}}(\cdot)=\lim_{p\to\infty}\|\eta_{+}^{\mathtt{t}}\|_{L^{p}(\nu(\cdot\cap\{\zeta^{\mathtt{t}}=0,\eta^{\mathtt{t}}\neq 0\}))}.$
Thus, $\mathtt{t}\mapsto\overline{\eta}^{\mathtt{t}}(0)$ is $\mathcal{T}$-measurable. ∎
As before, admissible strategies are signal-dependent investment choices $\tilde{\phi}$ which are measurable with respect to the Meyer-$\sigma$-field
$\Lambda:=\mathcal{P}(\bar{\mathcal{F}})\vee\sigma(Z).$
That is, we consider strategies $\tilde{\phi}:[0,T]\times\bar{\Omega}\to\mathbb{R}$ which by Lemma 2.1 (applied with $\bar{\Omega}$ instead of $\Omega$) can be written as
(29) $\tilde{\phi}_{t}(\bar{\omega})=\hat{\phi}_{t}(\bar{\omega},Z_{t}(\bar{\omega}))=\hat{\phi}^{\mathtt{t}}_{t}(\omega,Z^{\mathtt{t}}_{t}(\omega)),\quad(\bar{\omega},t)=(\mathtt{t},\omega,t)\in\bar{\Omega}\times[0,T]$
for some $\mathcal{P}(\bar{\mathcal{F}})\otimes\mathcal{B}(\mathbb{R})$-measurable field $\hat{\phi}:[0,T]\times\bar{\Omega}\times\mathbb{R}\to\mathbb{R}$. The set of admissible strategies is
[…]
Although $\bar{\mathcal{A}}$ is very similar to the definition in the multi-agent game, here we insist on $\bar{\mathcal{A}}$ being a set of positions for the representative investor, not for an investor of a specific type.
Proposition 3.3.
Given an admissible strategy $\tilde{\phi}_{t}=\hat{\phi}_{t}(Z_{t})\in\bar{\mathcal{A}}$, the representative investor’s wealth process $X^{\tilde{\phi}}:=X^{\hat{\phi}}$ has dynamics
(30) $\begin{split}X^{\hat{\phi}}_{0}&=x_{0}^{\tau}\\ \frac{dX^{\hat{\phi}}_{t}}{X^{\hat{\phi}}_{t-}}&=(r^{\tau}+\hat{\phi}_{t}(0)(\kappa^{\tau}-r^{\tau}))dt+\sigma^{\tau}\hat{\phi}_{t}(0)dW_{t}+\sigma^{0\tau}\hat{\phi}_{t}(0)dW^{0}_{t}+\int_{E}\eta^{\tau}(e)\hat{\phi}_{t}(\zeta^{\tau}(e))N(dt,de),\end{split}$
for $t\in[0,T]$ and it remains uniformly bounded away from zero on $[0,T]$$\bar{\mathbb{P}}$-almost surely.
The proof of this uniform boundedness is an immediate consequence of the following result:
Lemma 3.4.
Independently from the strategy applied, the returns from jumps are uniformly bounded from below: There exists a constant $c>0$ such that for all $\tilde{\phi}=\hat{\phi}(Z)\in\bar{\mathcal{A}}$:
$0<c\leq(\mathbb{T}(d\mathtt{t})\otimes\nu(de))\text{-}\operatorname*{ess\,inf}(1+\phi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e))\eta^{\mathtt{t}}(e)).$
Proof.
If $|\operatorname*{supp}\mathbb{T}|<\infty$, we easily argue as in Lemma 2.5. If $|\operatorname*{supp}\mathbb{T}|=\infty$, we have by our choice of admissible positions (28) in that case that
$1+\phi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e))\eta^{\mathtt{t}}(e)\geq 1+\min\left\{-\frac{1-a}{\underline{\eta}^{\mathtt{t}}(\zeta^{\mathtt{t}}(e))}\cdot\underline{\eta}^{\mathtt{t}}(\zeta^{\mathtt{t}}(e)),\ -\frac{1-b}{\overline{\eta}^{\mathtt{t}}(\zeta^{\mathtt{t}}(e)))}\cdot\bar{\eta}^{\mathtt{t}}(\zeta^{\mathtt{t}}(e))\right\}=\min\{a,b\}>0$
for all $e\in E$ and all $\mathtt{t}\in\mathtt{T}$. ∎
As in the multi-agent game, we will prove existence of equilibria for the special class of signal-driven strategies:
Definition 3.5.
An admissible strategy $\tilde{\phi}=\hat{\phi}(Z)\in\bar{\mathcal{A}}$ is called (purely) signal-driven if $\hat{\phi}$ in (29) does not depend on $t$ and $\omega$, that is, if it can be written in the form
(31) $\tilde{\phi}_{t}(\bar{\omega})=\hat{\phi}_{t}(\bar{\omega},z)=\phi^{\mathtt{t}}(z),\quad(\bar{\omega},t)\in\bar{\Omega}\times[0,T]$
for a $\mathcal{T}\otimes\mathcal{B}(\mathbb{R})$-measurable time-independent function $\phi:\mathtt{T}\times\mathbb{R}\to\mathbb{R},\ (\mathtt{t},z)\mapsto\phi^{\mathtt{t}}(z)$. We denote by
[…]
the set of admissible signal-driven strategies.
We will limit our search for (yet to be defined) mean field equilibria to signal-driven ones, following the same reasoning as in the multi-agent case. Also, we will, in preparation for our proof of existence of an equilibrium, solve a best response problem for the representative investor given a fixed environment. This environment is again described through the geometric mean wealth as induced by a signal-driven strategy $\pi=\pi^{\mathtt{t}}\in\bar{\mathcal{A}}_{\text{sig}}$ which specifies each type’s $\mathtt{t}$ signal-driven strategy $\pi^{\mathtt{t}}$. Indeeed, denoting by $\bar{\mathbb{E}}[\cdot]$ the expected value over the product space $\bar{\Omega}$ with respect to $\bar{\mathbb{P}}$, we define the average wealth in the mean field setting as
$\bar{X}^{\pi}_{t}:=\exp\bar{\mathbb{E}}[\log X^{\pi}_{t}|\mathcal{F}^{0}_{T}].$
In light of our considerations for the limiting average wealth with infinitely many agents in Section 2.5, our next result shows that this is consistent with our initial derivation of average wealth in Lemma 2.12:
Lemma 3.6.
For any $\pi\in\bar{\mathcal{A}}_{\text{sig}}$ and respective wealth process $X^{\pi}$ from Proposition 3.3, we find that $\bar{X}^{\pi}_{t}=\exp\bar{\mathbb{E}}[\log X^{\pi}_{t}|\mathcal{F}^{0}_{T}]$ is given by
(32) $\begin{split}\bar{X}^{\pi}_{t}=\bar{x}_{0}\exp\bigg{\{}\overline{\tau\pi}t+\overline{\sigma^{0}\pi}W^{0}_{t}+\int_{[0,t]\times E^{\mathtt{c}}}\log(\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}}))N(ds,de^{\mathtt{c}},E^{\mathtt{i}})\bigg{\}}\end{split}$
for $\bar{x}_{0}$, $\overline{\tau\pi}$, $\overline{\sigma^{0}\pi}$ as defined in (24) and $\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}})$ as defined in (25). In particular, the dynamics coincide with the limit from Lemma 2.12.
Proof.
Let $\pi=\pi^{\tau}\in\bar{\mathcal{A}}_{\text{sig}}$. We note that
$\bar{X}^{\pi}_{t}=\exp\bar{\mathbb{E}}[\log X^{\pi}_{t}|\mathcal{F}^{0}_{T}]=\exp\bar{\mathbb{E}}[\log X^{\pi}_{t}|\mathcal{F}^{0}_{t}],$
where the second equality holds because $X^{\pi}_{t}$ is independent of $(N(ds,de))_{s\in[t,T]}$ and $(W^{0}_{s}-W^{0}_{t})_{s\in[t,T]}$. By Itō’s formula, we easily find that
[…]
where, since $\pi^{\tau}$ is signal-driven, the strategies do not depend on $t$ and $\omega$ but on the signal. In particular, $\pi^{\tau}(z)$ is $\sigma(\tau)$-measurable for $z\in\mathbb{R}$. This is important, as in the following, we take conditional expectation with respect to the common noise $\mathcal{F}^{0}_{t}$ and then apply the exponential:
First, for the initial value, we find that
$\bar{x}_{0}:=\bar{X}^{\pi}_{0}=\exp\bar{\mathbb{E}}[\log x_{0}^{\tau}|\mathcal{F}^{0}_{t}]=\exp\int_{\mathtt{T}}\log(x_{0}^{\mathtt{t}})\mathbb{T}(d\mathtt{t}),$
because $x_{0}^{\tau}$, being $\sigma(\tau)$-measurable, is independent from the common noise. Similarly, for the drift term, we exploit that the lifted common noise drivers $W^{0}$ and $N$ actually only depend on $\omega$, while the market parameters and $\pi^{\tau}(0)$ only depend on the type $\mathtt{t}$, thus, the $ds$-integrand averages to
$\bar{\mathbb{E}}$ $\left[r+\pi^{\tau}(0)(\kappa-r)-\frac{1}{2}(\sigma^{2}+(\sigma^{0})^{2})(\pi^{\tau}(0))^{2}\big{|}\mathcal{F}^{0}_{t}\right]$
$\quad=\int_{\mathtt{T}}r^{\mathtt{t}}+\pi^{\mathtt{t}}(0)(\kappa^{\mathtt{t}}-r^{\mathtt{t}})-\frac{1}{2}((\sigma^{\mathtt{t}})^{2}+(\sigma^{0\mathtt{t}})^{2})(\pi^{\mathtt{t}}(0))^{2}\mathbb{T}(d\mathtt{t})=\overline{\tau\pi}.$
Further, since $W$, $\sigma^{0}$ and $\pi^{\tau}(0)$ are independent from the common noise drivers, we find that when taking conditional expectation with respect to $\mathcal{F}^{0}_{t}$ the $dW$-term vanishes as a martingale with zero expectation. However, for the $dW^{0}$-term, randomness remains in the types of the integrand, that is
[…]
for
$\bar{\mathbb{E}}[\sigma^{0}\pi^{\tau}(0)|\mathcal{F}^{0}_{t}]=\int_{\mathtt{T}}\sigma^{0\mathtt{t}}\pi^{\mathtt{t}}(0)\mathbb{T}(d\mathtt{t})=\overline{\sigma^{0}\pi}.$
Finally, for the Poisson integral, we have to average over the types and the idiosyncratic jump marks, that is
$\bar{\mathbb{E}}\bigg{[}\int_{[0,t]\times E}$ $\log\big{(}1+\pi^{\tau}(\zeta^{\tau}(e))\eta^{\tau}(e)\big{)}N(ds,de)\bigg{|}\mathcal{F}^{0}_{t}\bigg{]}=\bar{\mathbb{E}}\left[\sum_{t_{k}\leq t}\log\big{(}1+\pi^{\tau}(\zeta^{\tau}(e^{\mathtt{c}}_{k},e^{\mathtt{i}}_{k}))\eta^{\tau}(e^{\mathtt{c}}_{k},e^{\mathtt{i}}_{k})\big{)}\bigg{|}(t_{k})_{k},(e^{\mathtt{c}}_{k})_{k}\right]$
$=\sum_{t_{k}\leq t}\int_{E^{\mathtt{i}}\times\mathtt{T}}\log\big{(}1+\pi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e^{\mathtt{c}}_{k},e^{\mathtt{i}}))\eta^{\mathtt{t}}(e^{\mathtt{c}}_{k},e^{\mathtt{i}})\big{)}\nu^{i}(de^{\mathtt{i}})\otimes\mathbb{T}(d\mathtt{t})$
$=\int_{[0,t]\times E^{\mathtt{c}}}\bigg{(}\int_{E^{\mathtt{i}}\times\mathtt{T}}\log\big{(}1+\pi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}})\big{)}\nu^{i}(de^{\mathtt{i}})\otimes\mathbb{T}(d\mathtt{t})\bigg{)}N(ds,de^{\mathtt{c}},E^{\mathtt{i}})$
for the mark times $t_{k}$ and common and idiosyncratic jump marks $(e^{\mathtt{c}}_{k},e^{\mathtt{i}}_{k})\in E^{\mathtt{c}}\times E^{\mathtt{i}}$, $k=1,2,\dots$, set by the Poisson random measure $N$. Together with (25) and taking the exponential, we find (32) which coincides with the limit from Lemma 2.12. ∎
As we will see soon, the above geometric mean field wealth works as a sufficient statistics of the control’s environment enabling us to bypass the so-called master equation when considering best responses.
Let us fix some $\pi\in\bar{\mathcal{A}}_{\text{sig}}$ and resulting geometric mean wealth $\bar{X}^{\pi}_{T}=\exp\bar{\mathbb{E}}[\log X_{T}^{\pi}|\mathcal{F}^{0}_{T}]$. Then, the representative investor’s best response problem is to
[…]
subject to her wealth process $X^{\tilde{\phi}}:=X^{\hat{\phi}}$ from (30). Here, $u=u_{\tau}$ is the utility function dependent on the type $\tau$, which almost surely behaves as the utility defined in (6), i.e.
$u(\bar{\omega},x,\bar{x}):=u_{\tau(\bar{\omega})}(x,\overline{x})=u_{\mathtt{t}}(x,\overline{x})=u_{\alpha^{\mathtt{t}}}(x\bar{x}^{-\theta^{\mathtt{t}}})\text{ for }\bar{\omega}=(\mathtt{t},\omega)\in\mathtt{T}\times\Omega=\bar{\Omega}.$
Ultimately, an equilibrium strategy $\tilde{\pi}^{\star}\in\bar{\mathcal{A}}$ aligns the best response control with the control of the fixed environment. For this, Lemma 3.6 justifies the so-called consistency condition to be
$\bar{X}^{\tilde{\pi}^{\star}}_{T}=\exp\bar{\mathbb{E}}[\log X^{\tilde{\pi}^{\star}}_{T}|\mathcal{F}^{0}_{T}],$
and it, indeed, shows up in the definition of the mean field equilibrium:
Definition 3.7.
Let $\tilde{\pi}^{\star}=\hat{\pi}^{\star}(Z)\in\bar{\mathcal{A}}$ be an admissible strategy and consider $\bar{X}^{\tilde{\pi}^{\star}}_{T}=\exp\bar{\mathbb{E}}[\log X^{\tilde{\pi}^{\star}}_{T}|\mathcal{F}^{0}_{T}]$, where $X^{\tilde{\pi}^{\star}}$ is the wealth process corresponding to $\tilde{\pi}^{\star}$. We say that $\tilde{\pi}^{\star}$ is a mean field equilibrium if
(33) $\bar{\mathbb{E}}[u(X^{\tilde{\pi}}_{T},\bar{X}_{T}^{\tilde{\pi}^{\star}})]\leq\bar{\mathbb{E}}[u(X^{\tilde{\pi}^{\star}}_{T},\bar{X}^{\tilde{\pi}^{\star}}_{T})]\text{ for all }\tilde{\pi}\in\bar{\mathcal{A}}.$
A mean field equilibrium $\pi^{\star}$ is called signal-driven if $\pi^{\star}\in\bar{\mathcal{A}}_{\text{sig}}$ is signal-driven in the sense of Definition 3.5.
As mentioned, we will will first solve the best response problem of the representative investor.
3.2. HJB equation and Best Response Map for The Mean Field Game
Let $\pi\in\bar{\mathcal{A}}_{\text{sig}}$ describe the investment decisions made in the mean field and consider the resulting geometric mean wealth $\bar{X}^{\pi}_{T}=\exp\bar{\mathbb{E}}[\log X_{T}^{\pi}|\mathcal{F}^{0}_{T}]$ from (32).
The following lemma collects some technical results ensuring, in particular, that the mean wealth behaves ‘nicely’.
Lemma 3.8.
We find that
  1. (i)
    signal-driven investment positions are uniformly bounded: There exists a constant $C>0$ such that for all $\pi\in\bar{\mathcal{A}}_{\text{sig}}$ and for $\mathbb{T}$-almost every $\mathtt{t}\in\mathtt{T}$
    $\max_{z\in\mathbb{R}}|\pi^{\mathtt{t}}(z)|\leq C.$
  2. (ii)
    there exists a constant $\bar{c}>0$ such that $0<\bar{c}\leq\nu^{\mathtt{c}}(de^{\mathtt{c}})\text{-}\operatorname*{ess\,inf}\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}})$ for all $\pi\in\bar{\mathcal{A}}_{\text{sig}}$,
  3. (iii)
    for all $p>0$ there exists a $L^{p}(\nu^{c})$-integrable function $f:E^{c}\to\mathbb{R}$ such that for all $\pi\in\bar{\mathcal{A}}_{\text{sig}}$, we have
    $\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}})\leq f(e^{c}).$
Further, from (ii) and (iii) it follows that for $\mathbb{T}$-almost every $\mathtt{t}\in\mathtt{T}$, for all $z\in\mathbb{R}$ and any $\varphi\in\Phi^{\mathtt{t}}(z)$ there exists a $\nu$-integrable function $g:E\to\mathbb{R}$ such that for all $\pi\in\bar{\mathcal{A}}_{\text{sig}}$ we have
(iv) $\begin{split}\left|u_{\mathtt{t}}(1+\varphi\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}),\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}}))\right|\leq g(e).\end{split}$
Note that, by (ii), the mean wealth dynamics in (32) remain strictly nonnegative. Also, (i) together with $(\mathtt{T}_{4})$, ensures that $\overline{\tau\pi}$ is finite and $\sigma^{0}\pi(0)$ is uniformly bounded with respect to $\mathbb{T}$ independent from the choice of $\pi$. Further, (iii) and (iv) ensure finiteness of the suprema of the HJB equation below, see (37), as well as sufficient continuity of the target funtions of the best response maps, see (40) and (41), and, later, also continuity of the fixed point maps in the proof of existence of mean field equilibria.
Proof.
Let $\pi\in\bar{\mathcal{A}}_{\text{sig}}$. For (i), it is straightforward that ($\mathtt{T}_{5}$) bounds signal-driven investment positions uniformly in types and signals independently from $\pi$ as
$\max_{z\in\mathbb{R}}|\pi^{\mathtt{t}}(z)|\leq\max_{z\in\mathbb{R}}\max\left\{\bigg{|}\frac{1}{\bar{\eta}^{\mathtt{t}}(z)}\bigg{|},\bigg{|}\frac{1}{\underline{\eta}^{\mathtt{t}}(z)}\bigg{|}\right\}\leq|(c_{0})^{-1}|.$
For (ii), it suffices to show that there exists $\tilde{c}\in\mathbb{R}$ independent from $\pi$ such that for all $e^{\mathtt{c}}\in E^{\mathtt{c}}$, we have that
$-\infty<\tilde{c}\leq(\nu^{i}(de^{\mathtt{i}})\otimes\mathbb{T}(d\mathtt{t}))\text{-}\operatorname*{ess\,inf}\big{(}\log(1+\pi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\big{)},$
because then the exponential $\overline{1+(\pi\circ\zeta)\eta}$ is uniformly bounded away from zero on $E^{\mathtt{c}}$; and indeed, this follows immediately from Lemma 3.4.
For (iii), observe that for all $e^{\mathtt{c}}\in E^{\mathtt{c}}$, by (i), we find
$\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}})$ $\leq\int_{\mathtt{T}\times E^{\mathtt{i}}}\big{(}1+\pi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}})\big{)}\nu^{i}(de^{\mathtt{i}})\otimes\mathbb{T}(d\mathtt{t})$
$\leq(1+|(c_{0})^{-1}|)\int_{\mathtt{T}\times E^{\mathtt{i}}}(1+|\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}})|)\nu^{i}(de^{\mathtt{i}})\otimes\mathbb{T}(d\mathtt{t})$
where the right-hand side is independent of $\pi$ and implies existence of $f$ as above by $(\mathtt{T}_{6})$.
Lastly, concerning (iv), we find that
$\begin{split}\left|u_{\mathtt{t}}(1+\varphi\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}),\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}}))\right|\leq\frac{(1+\varphi\eta^{\mathtt{t}}(e))^{1-\alpha^{\mathtt{t}}}}{|1-\alpha^{\mathtt{t}}|}\big{(}\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}})\big{)}^{-\theta^{\mathtt{t}}(1-\alpha^{\mathtt{t}})}\end{split}$
where both for $\alpha^{\mathtt{t}}\in(0,1)$ and $\alpha^{\mathtt{t}}\in(1,\infty)$ the right-hand side exhibits a denominator which is uniformly bounded away from zero by Lemma 3.4 and (ii), and a nominator which, by $(\mathtt{T}_{6})$ together with (i) and (iii), is bounded from above by a $\nu$-integrable function which is independent from $\pi$. ∎
Returning to our initial goal, for given $\pi\in\bar{\mathcal{A}}_{\text{sig}}$, we search the best response $\tilde{\phi}^{\star}\in\bar{\mathcal{A}}$ determined by the problem to
(34) maximize $\text{maximize }\bar{\mathbb{E}}[u(X^{\tilde{\phi}}_{T},\bar{X}^{\pi}_{T})]\text{ over }\tilde{\phi}\in\bar{\mathcal{A}}.$
To make this problem amenable to dynamic programming we write
$\bar{\mathbb{E}}[u(X^{\tilde{\phi}}_{T},\bar{X}^{\pi}_{T})]=\int_{\mathtt{T}}\mathbb{E}[u_{\mathtt{t}}(X^{\tilde{\phi}^{\mathtt{t}}}_{T},\bar{X}^{\pi}_{T})]\mathbb{T}(d\mathtt{t})$
and introduce the value function for the representative investor’s problem when she is of type $\mathtt{t}\in\mathtt{T}$:
(35) $v^{\mathtt{t}}(x,\bar{x},T):=\sup_{\tilde{\phi}\in\bar{\mathcal{A}}^{\mathtt{t}}}\mathbb{E}[u_{\mathtt{t}}(X^{\tilde{\phi}}_{T},\bar{X}^{\pi}_{T})].$
Here we use the type-dependent set of admissible strategies
$\bar{\mathcal{A}}^{\mathtt{t}}:=\{\tilde{\phi}=\hat{\phi}(Z^{\mathtt{t}})\text{ for some $\mathcal{P}(\mathcal{F})\otimes\mathcal{B}(\mathbb{R})$-measurablefield $\hat{\phi}$ with }\hat{\phi}_{t}(z)\in\bar{\Phi}^{\mathtt{t}}(z)\text{for all }z\in\mathbb{R},\ t\in[0,T]\},$
where the filtration $\mathcal{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ on $\Omega$ is induced by the completions of
$\sigma((W^{0}_{s},W_{s},N([0,s]\times A)):s\in[0,t],A\in\mathcal{E}),\quad t\in[0,T],$
and for the resulting wealth dynamics:
(36) $\begin{split}X^{\hat{\phi}}_{0}&=x_{0}^{\mathtt{t}}\\ \frac{dX_{t}^{\hat{\phi}}}{X^{\hat{\phi}}_{t-}}&=(r^{\mathtt{t}}+\hat{\phi}_{t}(0)(\kappa^{\mathtt{t}}-r^{\mathtt{t}}))dt+\sigma^{\mathtt{t}}\hat{\phi}_{t}(0)dW_{t}+\sigma^{0\mathtt{t}}\hat{\phi}_{t}(0)dW^{0}_{t}+\int_{E}\hat{\phi}_{t}(\zeta^{\mathtt{t}}(e))\eta^{\mathtt{t}}(e)N(dt,de).\end{split}$
We will later also consider the type-dependent set of signal-driven admissible strategies which we define to be
$\bar{\mathcal{A}}^{\mathtt{t}}_{\text{sig}}:=\{\tilde{\phi}\in\bar{\mathcal{A}}^{\mathtt{t}}\text{ such that }\tilde{\phi}_{t}(\omega)=\phi(Z^{\mathtt{t}}_{t}(\omega))\text{ for }(\omega,t)\in\Omega\times[0,T]\text{ for some }\text{$\mathcal{B}(\mathbb{R})$-measurable }\phi:\mathbb{R}\to\mathbb{R}\}.$
Using the wealth dynamnics (32) and (36), we can follow the same heuristics as in the multi-agent game to derive the corresponding HJB equation:
(37) $\begin{split}&-\partial_{T}v^{\mathtt{t}}(x,\bar{x},T)+\partial_{x}v^{\mathtt{t}}(x,\bar{x},T)xr^{\mathtt{t}}+\partial_{y}v^{\mathtt{t}}(x,\bar{x},T)y\big{(}\overline{\tau\pi}-\tfrac{1}{2}(\overline{\sigma^{0}\pi})^{2}\big{)}+\tfrac{1}{2}\partial_{yy}v^{\mathtt{t}}(x,\bar{x},T)y^{2}(\overline{\sigma^{0}\pi})^{2}\\ &\quad+\sup_{\varphi\in\bar{\Phi}^{\mathtt{t}}(0)}\bigg{\{}\partial_{x}v^{\mathtt{t}}(x,\bar{x},T)x\varphi(\kappa^{\mathtt{t}}-r^{\mathtt{t}})+\tfrac{1}{2}\partial_{xx}v^{\mathtt{t}}(x,\bar{x},T)x^{2}\varphi^{2}((\sigma^{\mathtt{t}})^{2}+(\sigma^{0\mathtt{t}})^{2})+\partial_{xy}v^{\mathtt{t}}(x,\bar{x},T)xy\sigma^{0\mathtt{t}}\varphi\overline{\sigma^{0}\pi}\\ &\quad\quad\quad\quad\quad\quad\quad\quad+\int_{\{\zeta^{\mathtt{t}}=0\}}\bigg{[}v^{\mathtt{t}}(x(1+\varphi\eta^{\mathtt{t}}(e)),\bar{x}(\overline{1+(\pi\circ\zeta)\eta}(e)),T)-v^{\mathtt{t}}(x,\bar{x},T)\bigg{]}\nu(de)\bigg{\}}\\ &\quad+\int_{E\setminus\{0\}}\sup_{\varphi\in\bar{\Phi}^{\mathtt{t}}(z)}\bigg{\{}\int_{\{\zeta^{\mathtt{t}}=z\}}\bigg{[}v^{\mathtt{t}}(x(1+\varphi\eta(e)),\bar{x}(\overline{1+(\pi\circ\zeta)\eta}(e)),T)-v^{\mathtt{t}}(x,\bar{x},T)\bigg{]}K^{\mathtt{t}}(z,de)\bigg{\}}\mu^{\mathtt{t}}(dz)=0.\end{split}$
together with the boundary condition
$v^{\mathtt{t}}(x,\bar{x},0)=u_{\mathtt{t}}(x,\bar{x})=u_{\alpha^{\mathtt{t}}}(x\bar{x}^{-\theta^{\mathtt{t}}}),$
and where, for simplicity of notation, we write $\overline{1+(\pi\circ\zeta)\eta}(e):=\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}})$ for $e=(e^{\mathtt{c}},e^{\mathtt{i}})\in E$, cf. (25). Moreover, we provide the following verification result:
Lemma 3.9.
Fix $\pi\in\bar{\mathcal{A}}_{\text{sig}}$. Then, for $\mathbb{T}$-almost every $\mathtt{t}$, the value function $v^{\mathtt{t}}$ of (35) is given by
(38) $v^{\mathtt{t}}(x,\bar{x},T)=u_{\alpha^{\mathtt{t}}}(x\bar{x}^{-\theta^{\mathtt{t}}})e^{T(1-\alpha^{\mathtt{t}})((1-\theta^{\mathtt{t}})r^{\mathtt{t}}+M^{\mathtt{t}}_{\pi})}\quad\text{ for }(t,x,\bar{x})\in[0,T]\times(0,\infty)\times(0,\infty)$
for the finite constant
(39) $\begin{split}M^{\mathtt{t}}_{\pi}&:=-\theta^{\mathtt{t}}\overline{\tau\pi}+\theta^{\mathtt{t}}(\overline{\sigma^{0}\pi})^{2}+\tfrac{1}{2}(\theta^{\mathtt{t}})^{2}(1-\alpha^{\mathtt{t}})(\overline{\sigma^{0}\pi})^{2}\\ &\quad\quad+\sup_{\varphi\in\bar{\Phi}^{\mathtt{t}}(0)}\bigg{\{}\varphi(\kappa^{\mathtt{t}}-r^{\mathtt{t}})-\tfrac{1}{2}\alpha^{\mathtt{t}}\varphi^{2}((\sigma^{\mathtt{t}})^{2}+(\sigma^{0\mathtt{t}})^{2})-\theta^{\mathtt{t}}(1-\alpha^{\mathtt{t}})\sigma^{0\mathtt{t}}\varphi\overline{\sigma^{0}\pi}\\ &\qquad\qquad\qquad\qquad\qquad\qquad+\int_{\{\zeta^{\mathtt{t}}=0\}}\bigg{[}u_{\mathtt{t}}(1+\varphi\eta^{\mathtt{t}}(e),\overline{1+(\pi\circ\zeta)\eta}(e))-u_{\mathtt{t}}(1,1)\bigg{]}\nu(de)\bigg{\}}\\ &\quad\quad+\int_{\zeta^{\mathtt{t}}(E)\setminus\{0\}}\sup_{\varphi\in\bar{\Phi}^{\mathtt{t}}(z)}\bigg{\{}\int_{\{\zeta^{\mathtt{t}}=z\}}\bigg{[}u_{\mathtt{t}}(1+\varphi\eta^{\mathtt{t}}(e),\overline{1+(\pi\circ\zeta)\eta}(e)])-u_{\mathtt{t}}(1,1)\bigg{]}K^{\mathtt{t}}(z,de)\bigg{\}}\mu^{\mathtt{t}}(dz).\end{split}$
Moreover, the value function $v^{\mathtt{t}}$ is indeed a smooth solution of the HJB equation (37) and the best response control is again a signal-driven strategy $\tilde{\phi}_{\pi}^{\mathtt{t},\star}\in\bar{\mathcal{A}}^{\mathtt{t}}_{\text{sig}}$: at times $t$ without a signal, i.e., when $z=Z^{\mathtt{t}}_{t}=0$, it is given by the default investment
(40) $\begin{split}\phi_{\pi}^{\mathtt{t},\star}(0)&=\operatorname*{arg\,max}_{\varphi\in\bar{\Phi}^{\mathtt{t}}(0)}\bigg{\{}\varphi(\kappa^{\mathtt{t}}-r^{\mathtt{t}})-\tfrac{1}{2}\alpha^{\mathtt{t}}\varphi^{2}((\sigma^{\mathtt{t}})^{2}+(\sigma^{0\mathtt{t}})^{2})-\theta^{\mathtt{t}}(1-\alpha^{\mathtt{t}})\sigma^{0\mathtt{t}}\varphi\overline{\sigma^{0}\pi}\\ &\qquad\qquad\qquad\qquad\qquad\qquad+\int_{\{\zeta^{\mathtt{t}}=0\}}\bigg{[}u_{\mathtt{t}}(1+\varphi\eta^{\mathtt{t}}(e),\overline{1+(\pi\circ\zeta)\eta}(e))-u_{\mathtt{t}}(1,1)\bigg{]}\nu(de)\bigg{\}},\end{split}$
and, at moments $t$ when receiving a signal $z=Z^{\mathtt{t}}_{t}\neq 0$, it is best to invest the fraction
(41) $\phi_{\pi}^{\mathtt{t},\star}(z)=\operatorname*{arg\,max}_{\varphi\in\bar{\Phi}^{\mathtt{t}}(z)}\bigg{\{}\int_{\{\zeta^{\mathtt{t}}=z\}}\bigg{[}u_{\mathtt{t}}(1+\varphi\eta^{\mathtt{t}}(e),\overline{1+(\pi\circ\zeta)\eta}(e))-u_{\mathtt{t}}(1,1)\bigg{]}K^{\mathtt{t}}(z,de)\bigg{\}}.$
This lemma closely resembles the corresponding result, Theorem 2.10, in the multi-agent game. Consequently, the proof requires only minor adaptations to account for the different mean wealth. For sake of brevity, we omit it here and confine ourselves to pointing the reader to our standing Assumption 3.1 and Lemma 3.8 which give us all the tools to argue exactly as in the proof of Theorem 2.10.
Having thus found the best responses of each type of agent, we can now piece these together to find the best response control of our representative agent and thus solve her best response problem (34):
Theorem 3.10.
The best response control $\phi^{\star}_{\pi}\in\bar{\mathcal{A}}$ to $\pi\in\bar{\mathcal{A}}_{\text{sig}}$ in the sense of (34) is an admissible signal-driven strategy $\phi^{\star}_{\pi}=\phi^{\tau,\star}_{\pi}\in\bar{\mathcal{A}}_{\text{sig}}$ with $(\phi^{\mathtt{t},\star}_{\pi})_{\mathtt{t}\in\mathtt{T}}$ given by (40) and (41).
Proof.
For given $\pi\in\bar{\mathcal{A}}_{\text{sig}}$, we need to show that the strategy $\phi^{\star}_{\pi}=(\phi^{\mathtt{t},\star}_{\pi})_{\mathtt{t}\in\mathtt{T}}$ given by (40) and (41) maximizes
$\bar{\mathbb{E}}[u(X^{\tilde{\phi}}_{T},\bar{X}^{\pi}_{T})]$
over all $\tilde{\phi}$ on $\bar{\mathcal{A}}$. By Lemma 3.9, for any $\tilde{\phi}\in\bar{\mathcal{A}}$, it holds that
(42) $\bar{\mathbb{E}}[u(X^{\tilde{\phi}}_{T},\bar{X}^{\pi}_{T})]=\int_{\mathtt{T}}\mathbb{E}[u_{\mathtt{t}}(X^{\tilde{\phi}^{\mathtt{t}}}_{T},\bar{X}^{\pi}_{T})]\mathbb{T}(d\mathtt{t})\leq\int_{\mathtt{T}}\sup_{\tilde{\phi}^{\mathtt{t}}\in\bar{\mathcal{A}}^{\mathtt{t}}}\mathbb{E}[u_{\mathtt{t}}(X^{\tilde{\phi}^{\mathtt{t}}}_{T},\bar{X}^{\pi}_{T})]\mathbb{T}(d\mathtt{t})=\int_{\mathtt{T}}v^{\mathtt{t}}(x,y,0)\mathbb{T}(d\mathtt{t}).$
Defining $\phi^{\star}_{\pi}(\bar{\omega})=\phi^{\mathtt{t},\star}_{\pi}(Z^{\mathtt{t}}(\omega))$ for $\bar{\omega}=(\mathtt{t},\omega)\in\bar{\Omega}$ with $\phi^{\mathtt{t},\star}_{\pi}(z)$ given by (40) and (41), we find that $\phi^{\star}_{\pi}$ indeed attains the upper bound in (42) as soon as we have shown that $\phi^{\star}_{\pi}$ is a $\mathcal{T}\otimes\mathcal{B}(\mathbb{R})$-measurable field. Since, by $(\mathtt{T}_{3})$, there are only finitely many signals for each type, it sufficies to ensure measurable dependence of (40) and (41) on $\mathtt{t}\in\mathtt{T}$. For this, note that the target functions in (40) and (41) are Carathéodory functions: they are continuous in $\varphi$ by dominated convergence using Lemma 3.8 (iv), and as functions of $\mathtt{t}$ they are $\mathcal{T}$-measurable. Together with Lemma 3.2 this allows us to apply [AB06, Theorem 18.19] to show that the unique maximizers are indeed $\mathcal{T}$-measurable. In conclusion, $\phi^{\star}_{\pi}$ is an admissible signal-driven strategy and the best response control to $\pi$ in the sense of (34). ∎
3.3. Existence of a Mean Field Equilibrium
Similarly to the multi-agent game, the main task now is to define appropriate domain and image sets which ensure that Schauder’s Fixed Point Theorem can be applied to a suitable fixed point mapping. However, here, instead of finitely many investors, we will rely on a restriction to finitely many types of investors, or on a restriction to finitely many different common noise marks, i.e. $|E|^{\mathtt{c}}<\infty$. Consequently, we will provide two existence results, the first of which treats the setting of finitely many types:
Theorem 3.11.
Under Assumption 3.1, there exists a signal-driven mean field equilibrium $\pi^{\star}$ if $|\operatorname*{supp}\mathbb{T}|<\infty$.
Proof.
We need to find $\pi^{\star}\in\bar{\mathcal{A}}_{\text{sig}}$ which coincides with its best response, i.e. which satisfies $\phi^{\star}_{\pi^{\star}}=\pi^{\star}$, where $\phi^{\star}_{\pi^{\star}}$ is the optimal best control from Theorem 3.10 corresponding to $\pi^{\star}$. In fact, we, then, have that the consistency condition for the mean field equilibrium
$\bar{X}^{\pi^{\star}}_{T}=\exp\bar{\mathbb{E}}[\log X_{T}^{\pi^{\star}}|\mathcal{F}^{0}_{T}]=\exp\bar{\mathbb{E}}[\log X_{T}^{\phi^{\star}_{\pi^{\star}}}|\mathcal{F}_{T}^{0}],$
holds.
This fixed point problem can be reformulated as follows. Let us recall that $\operatorname*{supp}\mathbb{T}=:\bar{\mathtt{T}}\subset\mathtt{T}$ is finite by assumption and that each type can receive only finitely many distinct signals due to assumption $(\mathtt{T}_{3})$. Therefore, the set of admissible signal-driven strategies $\bar{\mathcal{A}}_{\text{sig}}$ can be identified with the compact, convex and non-empty set
$\bar{\Phi}:=\bigtimes_{\mathtt{t}\in\bar{\mathtt{T}}}\bigtimes_{z\in\zeta^{\mathtt{t}}(E)\cup\{0\}}\bar{\Phi}^{\mathtt{t}}(z).$
This is the domain and image set of the fixed-point mapping
$\gamma_{\text{MF}}:\bar{\Phi}\to\bar{\Phi},\quad\pi=(\pi^{\mathtt{t}})_{\mathtt{t}\in\bar{\mathtt{T}}}\mapsto\phi^{\star}_{\pi}=(\phi^{\mathtt{t},\star}_{\pi})_{\mathtt{t}\in\bar{\mathtt{T}}}.$
Let us show continuity of $\gamma_{\text{MF}}$. For this, first, we observe that, by Lemma 3.8(i) together with $(\mathtt{T}_{4})$ and Lemma 3.8(iii), both $\pi\mapsto\overline{\sigma^{0}\pi}$ and $\pi\mapsto\overline{1+(\pi\circ\zeta)}(e^{\mathtt{c}})$ are continuous on $\bar{\Phi}$, for each $e^{\mathtt{c}}\in E^{\mathtt{c}}$. Secondly, Lemma 3.8 (iv) allows us by dominated convergence to deduce that the target functions of the best response control $\phi^{\mathtt{t},\star}_{\pi}$ in (40) and (41) are continuous in $(\varphi,\pi)$. Thus, by Berge’s Maximum Theorem (cf. [AB06, Theorem 17.31]) the maximizers (40) and (41) are upper-hemicontinuous in $\pi$. Since these $\operatorname*{arg\,max}$-sets are actually singletons, the maximizers are continuously depending on $\pi$. As a consequence, $\gamma_{\text{MF}}$ is indeed continuous on $\bar{\Phi}$.
Finally, since $\bar{\Phi}$ is non-empty, compact and convex and $\gamma_{\text{MF}}$ is a continuous function, there exists $\pi^{\star}\in\bar{\Phi}$ such that $\pi^{\star}=\gamma_{\text{MF}}(\pi^{\star})$ by Schauder’s Fixed Point Theorem, cf. [Sch30]. ∎
Note that the previous proof uses the assumption $|\operatorname*{supp}\mathbb{T}|<\infty$ in an essential way to conclude continuity of the fixed-point mapping on a compact set. Indeed, when $|\operatorname*{supp}\mathbb{T}|=\infty$, defining an appropriate domain and image set for such a fixed-point mapping becomes a more delicate task as we can no longer rely on a finite-dimensional representation of the representative investor’s strategy. In the following we propose a way to circumvent this technical issue by assuming that there are only finitely many common noise marks. Indeed, under this assumption, the next lemma provides existence of a suitable fixed point mapping.
Lemma 3.12.
Suppose $|\operatorname*{supp}\mathbb{T}|=\infty$ and $|E^{\mathtt{c}}|<\infty$. Then, there exists a constant $\bar{C}>0$ such that for all $\pi\in\bar{\mathcal{A}}_{\text{sig}}$, $\mathbb{T}$-almost every $\mathtt{t}$ and $\nu$-almost every $e=(e^{\mathtt{c}},e^{\mathtt{i}})\in E$, we have
(43) $|\log(1+\pi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))|\leq\bar{C}.$
Moreover, there exist a non-empty convex and compact $K\subset\mathbb{R}\times\mathbb{R}^{|E^{c}|}$ with the following properties:
  • (i)
    For all $\pi\in\bar{\mathcal{A}}_{\text{sig}}$, we have
    $\mathtt{m}_{\pi}:=(\mathtt{m}_{\pi}(0),(\mathtt{m}_{\pi}(e^{\mathtt{c}}))_{e^{\mathtt{c}}\in E^{\mathtt{c}}}):=(\overline{\sigma^{0}\pi},(\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}}))_{e^{\mathtt{c}}\in E^{\mathtt{c}}})\in K.$
  • (ii)
    Every $m=(m(0),(m(e^{\mathtt{c}}))_{e^{\mathtt{c}}\in E^{\mathtt{c}}})\in K$, yields an admissible signal-driven strategy $\pi^{\star}_{m}\in\bar{\mathcal{A}}_{\text{sig}}$ where $\mathbb{T}$-a.e. type $\mathtt{t}$ will invest a fraction
    (44) $\begin{split}&\pi^{\mathtt{t},\star}_{m}(z)=\begin{cases}\operatorname*{arg\,max}\limits_{\varphi\in\bar{\Phi}^{\mathtt{t}}(0)}\bigg{\{}\varphi(\kappa^{\mathtt{t}}-r^{\mathtt{t}})-\tfrac{1}{2}\alpha^{\mathtt{t}}\varphi^{2}((\sigma^{\mathtt{t}})^{2}+(\sigma^{0\mathtt{t}})^{2})-\theta^{\mathtt{t}}(1-\alpha^{\mathtt{t}})\sigma^{0\mathtt{t}}\varphi m(0)&\\ \qquad\qquad\qquad\qquad+\displaystyle\int_{\{\zeta^{\mathtt{t}}=0\}}\bigg{[}u_{\mathtt{t}}(1+\varphi\eta^{\mathtt{t}}(e),m(e^{\mathtt{c}}))-u_{\mathtt{t}}(1,1)\bigg{]}\nu(de)\bigg{\}}&\text{ if }z=0,\\ \operatorname*{arg\,max}\limits_{\varphi\in\bar{\Phi}^{\mathtt{t}}(z)}\bigg{\{}\displaystyle\int_{\{\zeta^{\mathtt{t}}=z\}}\bigg{[}u_{\mathtt{t}}(1+\varphi\eta^{\mathtt{t}}(e),m_{\pi}(e^{\mathtt{c}}))-u_{\mathtt{t}}(1,1)\bigg{]}K^{\mathtt{t}}(z,de)\bigg{\}}&\text{ if }z\neq 0,\end{cases}\end{split}$
    of wealth in the stock when observing a signal $z$.
In particular,
(45) $\gamma_{\text{\rm MF}}(m):=\mathtt{m}_{\pi^{\star}_{m}}$
yields a mapping $\gamma_{\text{\rm MF}}:K\to K$.
Proof.
First, we show (43). For this, let $\pi\in\bar{\mathcal{A}}_{\text{sig}}$, $\mathtt{t}\in\mathtt{T}$ and $e=(e^{\mathtt{c}},e^{\mathtt{i}})\in E$. We find that, on $\{0<1+\pi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}})<1\}$, the $\log$-term is negative but uniformly bounded from below by Lemma 3.4. For the complementary case, i.e., on $\{1<1+\pi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}})\}$, the $\log$-term is positive and we find
$|\log(1+\pi^{\mathtt{t}}$ $(\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))|\leq\big{(}\max_{z\in\zeta^{\mathtt{t}}(E)\cup\{0\}}|\pi^{\mathtt{t}}(z)|\big{)}(1+\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}})),$
where both factors on the right-hand side are uniformly bounded by Lemma 3.8 (i) and $(\mathtt{T}_{6})$ (read for the present case $|\operatorname*{supp}\mathbb{T}|=\infty$).
For the the existence of $K$, we have that $\sigma^{0}\pi(0)$ is bounded on $\mathtt{T}$ independently from the choice of $\pi\in\bar{\mathcal{A}}_{\text{sig}}$ by Lemma 3.8 (i) together with $(\mathtt{T}_{4})$. Also, for the mean jumps, Lemma 3.8(ii) provides the lower and (43) the upper bound, that is, we immediately, find that $\bar{c}\leq m_{\pi}(e^{\mathtt{c}})\leq\bar{C}$ for all $\pi\in\bar{\mathcal{A}}_{\text{sig}}$ for all $e^{\mathtt{c}}\in E^{\mathtt{c}}$. Putting everything together, this ensures existence of $K$.
Also, following the same arguments as in the proof of Theorem 3.10, it is straightforward to show that $\pi^{\star}_{m}$ as defined in (44) is admissible and signal-driven. ∎
Now, we are ready to prove the existence of a mean field equilibrium when there are not necessarily finitely many types.
Theorem 3.13.
Under Assumption 3.1, there exists a signal-driven mean field equilibrium $\pi^{\star}$ if $|\operatorname*{supp}\mathbb{T}|=\infty$ and $|E^{\mathtt{c}}|<\infty$.
Proof.
Let us show that the self-map $\gamma_{\text{MF}}$ from (45) is continuous on $K$. First, we show continuity of
$(m(0),(m(e^{\mathtt{c}}))_{e^{\mathtt{c}}\in E^{\mathtt{c}}})\mapsto\pi^{\mathtt{t},\star}(z)$
for all $z\in\zeta^{\mathtt{t}}(E)\cup\{0\}$, for $\mathbb{T}$-almost every $\mathtt{t}$. By Lemma 3.8 (iv) and dominated convergence, the target functions of the argmax-operators of $\pi^{\mathtt{t},\star}(z)$ from (44) are continuous in $(\varphi,m(0),m(e^{\mathtt{c}}))_{e^{\mathtt{c}}\in E^{\mathtt{c}}}$. Then, by Berge’s Maximum Theorem (cf. [AB06, Theorem 17.31]) the $\operatorname*{arg\,max}$-operators are upper-hemicontinuous in $(m(0),(m(e^{\mathtt{c}}))_{e^{\mathtt{c}}\in E^{\mathtt{c}}})$. As these $\operatorname*{arg\,max}$-sets are actually singletons, the maximizers $\pi^{\mathtt{t},\star}_{m}(z)$ in (44) are given by functions which are continuous in $(m(0),(m(e^{\mathtt{c}}))_{e^{\mathtt{c}}\in E^{\mathtt{c}}})$, for all $z\in\zeta^{\mathtt{t}}(E)\cup\{0\}$ and for $\mathbb{T}$-almost every $\mathtt{t}$.
Secondly, recalling that
(46) $\begin{split}\mathtt{m}_{\pi^{\star}_{m}}(0)&=\overline{\sigma^{0}\pi^{\star}_{m}}=\int_{\mathtt{T}}\sigma^{0\mathtt{t}}\pi^{\mathtt{t},\star}_{m}(0)\mathbb{T}(d\mathtt{t}),\\ \mathtt{m}_{\pi^{\star}_{m}}(e^{\mathtt{c}})&=\overline{1+(\pi^{\star}_{m}\circ\zeta)}(e^{\mathtt{c}})=\exp\left\{\int_{\mathtt{T}\times E^{\mathtt{i}}}\log(1+\pi^{\mathtt{t},\star}_{m}(\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\nu^{\mathtt{i}}(de^{\mathtt{i}})\otimes\mathbb{T}(d\mathtt{t})\right\}.\end{split}$
for all $e^{\mathtt{c}}\in E^{\mathtt{c}}$, we find that the continuous dependence on $m=(m(0),(m(e^{\mathtt{c}}))_{e^{\mathtt{c}}\in E^{\mathtt{c}}})$ carries over because both integrands are uniformly bounded by Lemma 3.8(i), $(\mathtt{T}_{4})$ and (43), allowing us to apply dominated convergence. Thus, $\gamma_{\text{MF}}$ is continuous on $K$.
As a consequence, $\gamma_{\text{MF}}$ is a continuous self-map on $K$. Since $K\subset\mathbb{R}\times\mathbb{R}^{|E^{\mathtt{c}}|}$ is a non-empty compact and convex set, Schauder’s Fixed Point Theorem [Sch30] gives existence of a fixed point for $\gamma_{\text{MF}}$. ∎
Remark 3.14 (Uniqueness of Equilibria).
The uniqueness of the found equilibria in the above multi-agent and mean-field games remains an open question. In [LZ19], the authors prove the uniqueness within the class of constant equilibria for which they can rely on explicit best response controls in this very small class. We, however, consider merely signal-wise constant equilibrium strategies, for which the best response controls are given by one-dimensional concave optimization problems. Here, even in case of submodular utility, i.e. with $\alpha\in(0,1)$, uniqueness of the found equilibria seems far from straight-forward to prove, at leas to us.
This lack of uniqueness in equilibrium strategies prevents us from making any convergence statements based on these strategies and their aggregate wealth. Nevertheless, we note that our numerical experiments suggest uniqueness of the identified equilibrium. Indeed, in all our simulations, the fixed-point iteration always converged to the same equilibrium, irrespective of the randomly chosen strategies by which we started the iteration.
4. Case Studies and Financial-Economic Discussion
Equilibrium strategies are not available in explicit form, making equilibrium behavior difficult to characterize. Therefore, we turn to case studies and numerical methods to analyze how strategic interactions shape investor welfare and risk perception and to address questions like how much we should care about what others know.
In the following case studies, we will examine scenarios where investors differ in signal quantity, signal quality, and competitiveness. By systematically varying these parameters, we will identify how they impact the certainty equivalent assessing investors happiness and perceived risk. Our findings reveal when and why investors prefer certain competitive dynamics, whether they benefit from informational advantages or even favor better informed peers.
Our discussion focuses exclusively on mean field games for three reasons. First, we prefer scenarios where the influence of an individual on relative wealth is negligible and interactions are between the different types of players within the population. Secondly, and consequently, when individual decisions do not affect others, mean field games yield a more compelling than multi-agent games for our framework with its focus on (comparably simple) time-independent, signal-driven strategies. Finally, mean field games significantly reduce dimensionalty, simplifying the analysis and interpretation of investor behavior in large populations.
4.1. A Single Stock Model with Two Investor Types Receiving Different Signals
For simplicity, we consider mean field games with two different types $\mathtt{t}\in\{\mathtt{A},\mathtt{B}\}=:\mathtt{T}$ of investors trading in a single common stock subject to log-normal price shocks, that is, the parameters $\sigma=0$ and $r,\kappa,\sigma^{0},\eta$ do not vary with the type but are fixed. The proportions of types are given by $\mathbb{T}(\{\mathtt{A}\})=p^{\mathtt{A}}$, $\mathbb{T}(\{\mathtt{B}\})=p^{\mathtt{B}}\in[0,1]$ with $p^{\mathtt{A}}+p^{\mathtt{B}}=1$. We put $E^{\mathtt{c}}:=\mathbb{R}$ and $E^{\mathtt{i}}:=\mathbb{R}\times[0,1]$, and denote by $e^{\mathtt{i}}:=(e^{\mathtt{i}}_{1},e^{\mathtt{i}}_{2})\in E^{\mathtt{i}}$ a generic element in $E^{\mathtt{i}}$ which accounts for the independent idiosyncratic noise. Then, we can specify the jump intensity measures
$\nu^{\mathtt{c}}(de^{\mathtt{c}})=N_{0,1}(de^{\mathtt{c}})\text{ and }\nu^{\mathtt{i}}(de^{\mathtt{i}})=N_{0,1}(de_{1}^{\mathtt{i}})\otimes\mathcal{U}_{[0,1]}(de_{2}^{\mathtt{i}}),$
where $N_{0,1}$ denotes the standard normal distribution and $\mathcal{U}_{[0,1]}$ the uniform distribution on $[0,1]$. We define the common log-normal jumps as
$\eta(e):=\eta(e^{\mathtt{c}})=\exp(\hat{\sigma}e^{\mathtt{c}}+\hat{\kappa}-\tfrac{1}{2}\hat{\sigma}^{2})-1\quad\text{ for }e=(e^{\mathtt{c}},e^{\mathtt{i}})\in E:=E^{\mathtt{c}}\times E^{\mathtt{i}},$
for additional common market parameters $\hat{\kappa}\in\mathbb{R}$, $\hat{\sigma}>0$. Note that $\nu(\{\eta<-1\})=0$.
The idiosyncratic noise enters through the individual signals our agents receive. Each type of investor $\mathtt{t}\in\{\mathtt{A},\mathtt{B}\}$ receives, in case of an imminent jump in the stock price, a non-zero signal with an individual probability denoted by $p_{s}^{\mathtt{t}}\in[0,1]$. Such a non-zero signal will be a categorical indication about the sign and the size of a idiosyncratically noisy version of the impending common price shock $\eta(e)=\eta(e^{\mathtt{c}})$.
Specifically, we consider a perturbed version of this price price shock
$z^{\mathtt{t}}(e):=z^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}_{1}):=\rho^{\mathtt{t}}e^{\mathtt{c}}+\sqrt{1-(\rho^{\mathtt{t}})^{2}}e_{1}^{\mathtt{i}},$
where $|\rho^{\mathtt{t}}|<1$ determines the signal quality of an investor of type $\mathtt{t}\in\{\mathtt{A},\mathtt{B}\}$. Since, our existence theorems only work with finitely many distinct signals, we cannot use this $z^{\mathtt{t}}$ to define Type $\mathtt{t}$’s signal map $\zeta^{\mathtt{t}}$. Instead, we ensure $|\zeta^{\mathtt{t}}(E)|<\infty$ for $\mathtt{t}\in\{\mathtt{A},\mathtt{B}\}$ by categorizing the perturbed jump mark $z^{\mathtt{t}}(e)$ as “small”, “medium”, or “large” in absolute size and give an indication of the direction. Specifically, we put:
$\zeta^{\mathtt{t}}(e)=\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}_{1},e^{\mathtt{i}}_{2})=\mathds{1}_{[0,p_{s}^{\mathtt{t}}]}(e^{\mathtt{i}}_{2})\cdot\mathrm{sign}(z^{\mathtt{t}}(e))\cdot\begin{cases}0.5\quad&\text{ if }\quad|z^{\mathtt{t}}(e)|\leq 0.5,\\ 1\quad&\text{ if }\quad 0.5<|z^{\mathtt{t}}(e)|\leq 1,\\ +\infty\quad&\text{ if }\quad|z^{\mathtt{t}}(e)|>1,\end{cases}$
Thus, for both types the signal range is $\zeta^{\mathtt{t}}(E)\cup\{0\}=\{-\infty,-1,-0.5,0,0.5,1,\infty\}$. Still, signals may differ both in their frequency as parametrized by the probability $p^{\mathtt{t}}_{s}$ of signal reception and in their quality as determined by the perturbation parameter $\rho^{\mathtt{t}}$. Figure 1 and Figure 2 ilustrate this choice of jump signals.
Figure 1. For the original jump mark $e^{\mathtt{c}}=0.4$ with perturbed jump mark $z^{\mathtt{t}}(e)=0.8$ the signal $\zeta^{\mathtt{t}}(e)$ is $1$ indicating that the perturbed jump mark lies in the interval $(0.5,1]$.
Figure 2. For the original jump mark $e^{\mathtt{c}}=0.1$ with perturbed jump mark $z^{\mathtt{t}}(e)=-1.2$ the signal $\zeta^{\mathtt{t}}(e)$ is $\infty$- ∞ indicating that the perturbed jump mark lies in the interval $(-\infty,-1)$.
Specifying the following intervals of integration given a signal $z\in\zeta(E)$
$I(z):=\begin{cases}(1,+\infty)\quad&\text{ if }z=+\infty,\\ (0.5,1]\quad&\text{ if }z=1,\\ (0,0.5]\quad&\text{ if }z=0.5,\\ [-0.5,0)\quad&\text{ if }z=-0.5,\\ [-1,0.5)\quad&\text{ if }z=-1,\\ (-\infty,-1)\quad&\text{ if }z=-\infty,\\ \end{cases}$
we find the disintegration of $\nu(\cdot\cap\{\zeta^{\mathtt{t}}\neq 0\})$ for type $\mathtt{t}\in\{\mathtt{A},\mathtt{B}\}$ to result in the signal frequency
$\mu^{\mathtt{t}}(\{z\})$ $=\lambda p^{\mathtt{t}}_{s}N_{0,1}(I(z))\text{ for }z\in\zeta^{\mathtt{t}}(E)\setminus\{0\},$
and, for a signal $z\in\zeta^{\mathtt{t}}(E)\setminus\{0\}$, we find the a posteriori distribution for the jump marks to be
$K^{\mathtt{t}}(z,de)$ $=\frac{\mathds{1}_{I(z)}(z^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}_{1}))}{N_{0,1}(I(z))}N_{0,1}(de^{\mathtt{c}})\otimes N_{0,1}(de^{\mathtt{i}}_{1}).$
We impose that the probability of receiving a signal in case of a jump satisfies $p_{s}^{\mathtt{t}}<1$ and, hence, find that $\Phi^{\mathtt{t}}(z)=[0,1]$ for all $z\in\zeta(E)\cup\{0\}$.
In the remaining part of this paragraph, we explain how to efficiently compute the optimal investment positions for an investor of type $\mathtt{t}\in\{\mathtt{A},\mathtt{B}\}$. Readers may skip this section at first reading reading, and instead proceed directly to the numerical results and their discussion in Subsection 4.2.
For a signal-based strategy $\pi\in\bar{A}_{\text{sig}}$, we recall that the representative investor’s best reponse will be to invest a fraction
$\begin{split}\phi_{\pi}^{\mathtt{t},\star}(0)&=\operatorname*{arg\,max}_{\varphi\in\bar{\Phi}^{\mathtt{t}}(0)}\bigg{\{}\varphi(\kappa-r)-\tfrac{1}{2}\alpha^{\mathtt{t}}\varphi^{2}(\sigma^{2}+(\sigma^{0})^{2})-\theta^{\mathtt{t}}(1-\alpha^{\mathtt{t}})\sigma^{0}\varphi\overline{\sigma^{0}\pi}\\ &\qquad\qquad\qquad+\frac{\lambda(1-p_{s}^{\mathtt{t}})}{1-\alpha^{\mathtt{t}}}\bigg{[}\int_{\mathbb{R}}\big{[}(1+\varphi\eta(e^{\mathtt{c}}))^{1-\alpha^{\mathtt{t}}}(\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}}))^{-\theta^{\mathtt{t}}(1-\alpha^{\mathtt{t}})}-1\big{]}N_{0,1}(de^{\mathtt{c}})\bigg{]}\bigg{\}}\end{split}$
when not receiving a signal. We can explicitly compute
$\overline{\sigma^{0}\pi}=\bar{\mathbb{E}}[\sigma^{0}\pi(0)]=\sigma^{0}(p^{\mathtt{A}}\pi^{\mathtt{A}}(0)+p^{\mathtt{B}}\pi^{2}(0))$
and
(47) $\begin{split}&\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}})\\ &\quad=\exp\bigg{\{}\sum_{\mathtt{t}\in\{\mathtt{A},\mathtt{B}\}}p^{\mathtt{t}}\int_{E^{\mathtt{i}}}\log(1+\pi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\eta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}))\nu_{i}(de^{\mathtt{i}})\bigg{\}}\\ &\quad=\exp\bigg{\{}p^{\mathtt{A}}\bigg{(}(1-p^{\mathtt{A}}_{s})\log(1+\pi^{\mathtt{A}}(0)\eta(e^{\mathtt{c}}))+p^{\mathtt{A}}_{s}\int_{\mathbb{R}}\log(1+\pi^{\mathtt{A}}(\zeta^{\mathtt{A}}(e^{\mathtt{c}},e^{\mathtt{i}}_{1}))\eta(e^{\mathtt{c}}))N_{0,1}(de_{1}^{i})\bigg{)}\\ &\qquad\qquad\quad+p^{\mathtt{B}}\bigg{(}(1-p^{\mathtt{B}}_{s})\log(1+\pi^{\mathtt{B}}(0)\eta(e^{\mathtt{c}}))+p^{\mathtt{B}}_{s}\int_{\mathbb{R}}\log(1+\pi^{\mathtt{B}}(\zeta^{\mathtt{B}}(e^{\mathtt{c}},e^{\mathtt{i}}_{1}))\eta(e^{\mathtt{c}}))N_{0,1}(de_{1}^{\mathtt{i}})\bigg{)}\bigg{\}}.\end{split}$
Note that, given $e^{\mathtt{c}}\in E^{\mathtt{c}}$, a non-zero signal $0\neq z=\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}})$ allows one to deduce that the mark part $e^{\mathtt{i}}$ has to be in the interval
$I(z,e^{c}):=\begin{cases}\left(\frac{1-\rho^{\mathtt{t}}e^{\mathtt{c}}}{\sqrt{1-(\rho^{\mathtt{t}})^{2}}},+\infty\right)\quad&\text{ if }z=+\infty,\\ \left(\frac{z-0.5-\rho^{\mathtt{t}}e^{\mathtt{c}}}{\sqrt{1-(\rho^{\mathtt{t}})^{2}}},\frac{z-\rho^{\mathtt{t}}e^{\mathtt{c}}}{\sqrt{1-(\rho^{\mathtt{t}})^{2}}}\right)\quad&\text{ if }z=0.5\text{ or }z=1,\\ \left(\frac{z-\rho^{\mathtt{t}}e^{\mathtt{c}}}{\sqrt{1-(\rho^{\mathtt{t}})^{2}}},\frac{z+0.5-\rho^{\mathtt{t}}e^{\mathtt{c}}}{\sqrt{1-(\rho^{\mathtt{t}})^{2}}}\right)\quad&\text{ if }z=-0.5\text{ or }z=-1,\\ \left(-\infty,\frac{-1-\rho^{\mathtt{t}}e^{\mathtt{c}}}{\sqrt{1-(\rho^{\mathtt{t}})^{2}}}\right)\quad&\text{ if }z=-\infty,\\ \end{cases}$
Hence, the integrals in (47) simplify:
$\int_{\mathbb{R}}\log(1+\pi^{\mathtt{t}}(\zeta^{\mathtt{t}}(e^{\mathtt{c}},e^{\mathtt{i}}_{1}))\eta(e^{\mathtt{c}}))N_{0,1}(de_{1}^{\mathtt{i}})=\sum_{z\in\zeta^{\mathtt{t}}(E)\setminus\{0\}}\log(1+\pi^{\mathtt{t}}(z)\eta(e^{\mathtt{c}}))N_{0,1}(I(z,e^{c}))$
which is easily computed using the cummulative distribution function of the standard normal distribution.
Similarly, we can compute the components of the best response to $\pi$ when receiving a non-zero signal $z\neq 0$:
$\phi_{\pi}^{\mathtt{t},\star}(z)$ $=\operatorname*{arg\,max}_{\varphi\in\bar{\Phi}^{\mathtt{t}}(z)}\bigg{\{}\int_{\mathbb{R}}\bigg{[}\frac{(1+\varphi\eta(e^{\mathtt{c}}))^{1-\alpha^{\mathtt{t}}}(\overline{1+(\pi\circ\zeta)\eta}(e^{\mathtt{c}}))^{-\theta^{\mathtt{t}}(1-\alpha^{\mathtt{t}})}-1}{(1-\alpha^{\mathtt{t}})}\bigg{]}\frac{N_{0,1}(I(z,e^{c}))}{N_{0,1}(I(z))}N_{0,1}(de^{\mathtt{c}})\bigg{\}}.$
4.2. Numerical experiments
For all our experiments, we will consider the two types of investors to be of the same size in our population, i.e., $p^{\mathtt{A}}=p^{\mathtt{B}}=0.5$, and to have the same risk aversion $\alpha^{\mathtt{A}}=\alpha^{\mathtt{B}}=2$. Further, we choose the market parameters to be
$r=0,\ \kappa=0.08,\ \hat{\kappa}=0,\ \sigma^{0}=0.3,\hat{\sigma}=0.1,\lambda=10,$
and the time horizon for investments is $T=1$.
We investigate how the presence of investment signals impacts investors with relative performance concerns. For this, we consider investors of type $\mathtt{A}$ as a reference point and investigate how they assess their prospects when investors of type $\mathtt{B}$ are receiving more/less accurate or more/less frequent signals. To compare the outcomes of the different equilibria, let us introduce a suitable kind of certainty equivalent. The reference point for this certainty equivalent is in each case an equilibrium for the fully homogeneous reference type environment
$\mathtt{T}^{\text{ref}}$ =$=\{(x_{0}^{\mathtt{A},\text{ref}},p^{\mathtt{A},\text{ref}}_{s},\rho^{\mathtt{A},\text{ref}},\theta^{\mathtt{A},\text{ref}}),(x_{0}^{\mathtt{B},\text{ref}},p^{\mathtt{B},\text{ref}}_{s},\rho^{\mathtt{B},\text{ref}},\theta^{\mathtt{B},\text{ref}})\}$
=$=\{(1,0.5,0.5,0.5),(1,0.5,0.5,0.5)\}.$
We, then, let the characteristics of the investors of type $\mathtt{B}$ deviate from those in the reference scenario, that is, we consider an alternative type environment
$\mathtt{T}^{\text{alt}}$ =$=\{(x_{0}^{\mathtt{A},\text{ref}},p^{\mathtt{A},\text{ref}}_{s},\rho^{\mathtt{A},\text{ref}},\theta^{\mathtt{A},\text{ref}}),(x_{0}^{\mathtt{B},\text{alt}},p^{\mathtt{B},\text{alt}}_{s},\rho^{\mathtt{B},\text{alt}},\theta^{\mathtt{B},\text{alt}})\}$
=$=\{(1,0.5,0.5,0.5),(1,p^{\mathtt{B},\text{alt}}_{s},\rho^{\mathtt{B},\text{alt}},\theta^{\mathtt{B},\text{alt}})\},$
for various choices of $p^{\mathtt{B},\text{alt}}_{s}$, $\rho^{\mathtt{B},\text{alt}}$ and $\theta^{\mathtt{B},\text{alt}}$.
For both enviroments we compute the equilibrium strategies by numerical fixed point iterations. That is, we iteratively evaluate the best response controls from Section 4.1 using numerical integration and optimization methods until the iterations only yield negligible changes. The expected utilities for Type $\mathtt{A}$ investors will typically differ between these equilibria. We then determine the certainty equivalent as the initial capital $x_{0}^{\mathtt{A}}$ that an investor of type $\mathtt{A}$ would need in the reference setting $\mathtt{T}^{\text{ref}}$ to achieve the same expected utility there as in the alternative setting $\mathtt{T}^{\text{alt}}$:
$x_{0}^{\mathtt{A}}=\exp\left(M^{\mathtt{A},\text{alt}}_{\pi^{\star,\text{alt}}}-M^{\mathtt{A},\text{ref}}_{\pi^{\star,\text{ref}}}\right)$
for $M_{\pi^{\star,\text{ref}}}^{\mathtt{A},\text{ref}}$ and $M_{\pi^{\star,\text{alt}}}^{\mathtt{A},\text{alt}}$ as in (39) for corresponding type spaces $\mathtt{T}^{\text{ref}}$ and $\mathtt{T}^{\text{alt}}$ and equilibrium strategies $\pi^{\star,\text{ref}}$ and $\pi^{\star,\text{alt}}$, respectively.
Below, we present our results. For ease of presentation, we will sometimes refer to investors of type $\mathtt{A}$ as Type $\mathtt{A}$ and to investors of type $\mathtt{B}$ as Type $\mathtt{B}$. Also, whenever it is clear, we will drop the superscripts ‘ref’ and ‘alt’. Lastly, let us note that the lack of smoothness in some plots, notably in Figure 4, results from numerical approximation errors. The issue in these approximations arises from the computation of the integrals in the target functions of the best response strategies, which involve numerically delicate “spiky” integrands.
Better-Warned Peer Lowers Risk Exposures. Figure 3 describes how the certainty equivalent changes as the probability $p_{s}^{\mathtt{B},\text{alt}}$ for Type $\mathtt{B}$ to receive a signal increases, the other parameters being kept fixed their values in the reference setting. We find that the certainty equivalent increases with $p_{s}^{\mathtt{B},\text{alt}}$, indicating that investors of type $\mathtt{A}$ become happier when Type $\mathtt{B}$ is warned about price shocks more often.
More precisely, Type $\mathtt{A}$ investors find themselves confronted with competitive peers of Type $\mathtt{B}$ who, when warned less frequently, will have to take more risk in order to still catch up with the advantaged Type $\mathtt{A}$ investors. As a consequence, the average wealth $\bar{X}_{T}$ fluctuates more. These stronger fluctuations carry over to the relative performance of Type $\mathtt{A}$ investors whose risk aversion, then, makes them attribute a lower certainty equivalent value to such environments. Conversely, when Type $B$ is receiving more fequent signals, the added security that this affords also to Type $\mathtt{A}$ translates into higher certainty equivalents for the latter, even if it is harder for them to catch up with their peers under these circumstances.
Figure 3. Certainty equivalent $x_{0}^{\mathtt{A}}$ for equilibria where Type $\mathtt{B}$’s probability $p_{s}^{\mathtt{B},\text{alt}}$ of receiving a signal increaes from 0 to 1.
How much should we care about what others know? Figure 4 describes how Type $\mathtt{A}$ investors feel about their peers receiving more or less reliable signals when their signal frequency and concern for others is kept at the level in the reference equilibirium.
Figure 4. Certainty equivalent $x_{0}^{\mathtt{A}}$ for equilibria where Type $\mathtt{B}$’s signal quality $\rho^{\mathtt{B},\text{alt}}$ increases from 00 to $1$.
It reveals that they like a largely homogeneous signal quality, with the peak at $\rho^{\mathtt{B},\text{alt}}\approx 0.4$ revealing a slight preference for being blessed with a signal that is just a bit more reliable than that of the competition. When Type $\mathtt{B}$ peers receive more reliable signals, Type $\mathtt{A}$ finds it much harder to compete in wealth generation leading to a drop in the certainty equivalent for $\rho^{\mathtt{B},\text{alt}}>0.5$. The drop when $\rho^{\mathtt{B},\text{alt}}$ falls below $0.3$ is explained by Type $\mathtt{A}$’s risk aversion: here, Type $\mathtt{B}$ investors know that they have some catching up to do and thus take more risky positions, making also the relative performance of Type $\mathtt{A}$ investors more risky. When, however, their signal becomes too unreliable ($\rho^{\mathtt{B},\text{alt}}<0.2$) the risk aversion of type $\mathtt{B}$ investors makes them refrain from such aggressive positions in the stock market. This makes Type $\mathtt{A}$’s relative performance less volatile again, leading to the certainty equivalent’s increase close to $\rho^{\mathtt{B},\text{alt}}\approx 0$. In short: we should care a lot about the others’ knowledge if that knowledge is not as reliable as our own—and it does not matter whether our peers’ knowledge is significantly more or less reliable!
Furthermore, it is not just what others know that matters, but also how fiercely they act on that knowledge. The level of competitiveness among peers affects risk exposure and, consequently, the certainty equivalent. This dynamic is explored in the next two plots.
Increased Competitiveness Aligns Investor Objectives. Figure 5 shows that investors of type $\mathtt{A}$ always prefer an increase in the competitiveness of investors of type $\mathtt{B}$. This preference persists regardless of whether Type $\mathtt{B}$’s signal frequency is lower ($p_{s}^{\mathtt{B},\text{alt}}=0.1$), higher ($p_{s}^{\mathtt{B},\text{alt}}=0.9$), or at the same level ($p_{s}^{\mathtt{B},\text{alt}}=0.5$). This preference arises because the more competitive investors of type $\mathtt{B}$ are the more they strive to reduce fluctuations in relative mean wealth, which aligns with the objective of (also) risk-averse investors of Type $\mathtt{A}$.
Figure 5. Certainty equivalent $x_{0}^{\mathtt{A}}$ for equilibria as Type $\mathtt{B}$ concern for others $\theta^{\mathtt{B},\text{alt}}$ increases from 00 to $1$ for different levels of signal quantity $p_{s}^{\mathtt{B},\text{alt}}$.
Moreover, as in Figure 3, Type $\mathtt{A}$ investors prefer an environment where Type $\mathtt{B}$ is warned more frequently than they are, even if this means that they are operating at a disadvanage.
Additionally, the effect of increasing competitivness $\theta^{\mathtt{B},\text{alt}}$ on investors’ happiness is more pronounced when Type $\mathtt{B}$ is stronger in signal quantity. Since Figure 3 already shows that investors of Type $\mathtt{A}$ prefer better-warned peers, the combination of increased signal quantity and competitiveness further reinforces this preference. A better-informed and more competitive Type $\mathtt{B}$ takes proactive investment positions that reduce the fluctuations in relative wealth, very much to the pleasure of risk-averse Type $\mathtt{A}$ investors.
Preferences for Homogeneous over Heterogeneous Information. Figure 6 considers the same situation as Figure 5 but focusses on advantages and disadvantages in signal quality. Again, Type $\mathtt{A}$ prefers an increase in the competitiveness of Type $\mathtt{B}$. However, in contrast to Figure 5, Type $\mathtt{A}$ favors having an advantage in signal quality over being disadvantaged and, perhaps unexpectedly, prefers equal signal quality over holding an advantage. Knowing about their disadvantage, highly competitive Type $\mathtt{B}$ have strong incentives to compensate for their inferior signals by adopting more aggressive investment positions, increasing their wealth fluctuations and thus also the relative performance of their peers. In contrast, when signal quality is homogeneous, both investors react similarly to new information, fostering a more stable environment. As a result, Type $\mathtt{A}$ prefers equal signal quality over an advantage, not because an informational edge is inherently undesirable, but because the volatility created by a disadvantaged but aggressive competitor outweighs the benefits of superior information.
Figure 6. Certainty equivalent $x_{0}^{\mathtt{A}}$ for equilibria where the competitveness of Type $\mathtt{B}$ varies from $\theta^{\mathtt{B}}=0$ to $\theta^{\mathtt{B}}=1$ for different levels of signal quality $\rho^{\mathtt{B},\text{alt}}$.
Still, we note that the curves in Figure 6 are less steep compared to Figure 5, suggesting that the impact of increasing competitiveness is weaker when differences in signal quality, rather than quantity, drive the competition. This aligns with the earlier findings: disparities in signal quantity affect more directly how often investors adjust their positions, leading to larger shifts in market dynamics. While differences in signal quality introduce instability, they do so less forcefully, making competitiveness a less dominant factor in shaping investor preferences.